### v1

# LLM as GNN: Graph Vocabulary Learning for Text-Attributed Graph Foundation Models

Abstract
Text-Attributed Graphs (TAGs), where each node is associated with text descriptions, are ubiquitous in real-world scenarios. They typically exhibit distinctive structure and domain-specific knowledge, motivating the development of a Graph Foundation Model (GFM) that generalizes across diverse graphs and tasks. Despite large efforts to integrate Large Language Models (LLMs) and Graph Neural Networks (GNNs) for TAGs, existing approaches suffer from decoupled architectures with two-stage alignment, limiting their synergistic potential. Even worse, existing methods assign out-of-vocabulary (OOV) tokens to graph nodes, leading to graph-specific semantics, token explosion, and incompatibility with task-oriented prompt templates, which hinders cross-graph and cross-task transferability. To address these challenges, we propose PromptGFM, a versatile GFM for TAGs grounded in graph vocabulary learning. PromptGFM comprises two key components: (1) Graph Understanding Module, which explicitly prompts LLMs to replicate the finest GNN workflow within the text space, facilitating seamless GNN-LLM integration and elegant graph-text alignment; (2) Graph Inference Module, which establishes a language-based graph vocabulary ensuring expressiveness, transferability, and scalability, enabling readable instructions for LLM fine-tuning. Extensive experiments demonstrate our superiority and transferability across diverse graphs and tasks. The code is available at this URL 111 https://github.com/agiresearch/PromptGFM.
LLM as GNN: Graph Vocabulary Learning for Text-Attributed
Graph Foundation Models
Xi Zhu1thanks: Equal contribution., Haochen Xue211footnotemark: 1, Ziwei Zhao3, Wujiang Xu1, Jingyuan Huang1, Minghao Guo1, Qifan Wang4, Kaixiong Zhou5, Yongfeng Zhang1thanks: Corresponding email: yongfeng.zhang@rutgers.edu. 1 Rutgers University 2 University of Liverpool 3 University of Science and Technology of China 4 Meta AI 5 North Carolina State University
1 Introduction
Figure 1: Overview of three GNN-LLM integration paradigms for graph-text alignment: (a) GNN for LLM and (b) LLM for GNN use decoupled architectures in the embedding space, while (c) our work functions LLM as GNN in the text space.
Graphs, characterized by their non-Euclidean structures and rich domain-specific knowledge, serve as fundamental representations of complex relational data. Many of them integrate textual information at the node level, forming Text-Attributed Graphs (TAGs), such as citation networks  (Eto, 2019; Hu et al., 2020; Buneman et al., 2021), social networks  (Kempe et al., 2003; Myers et al., 2014), and molecular graphs  (Wieder et al., 2020; Jin et al., 2024a). However, existing solutions heavily depend on task- or dataset-specific training and deployment, limiting their transferability. To address this, we aim to build a Graph Foundation Model (GFM) capable of generalizing across different graphs and tasks for TAGs  (Mao et al., 2024; Xia et al., 2024).
In TAGs, Graph Neural Networks (GNNs) and Large Language Models (LLMs) are employed to handle different modalities, as illustrated in Figure 1. (a) GNN for LLM. GNNs generate structure-aware node embeddings that complement original textual embeddings, improving LLM inference  (Tang et al., 2024; Chai et al., 2023; Liu et al., 2024b). (b) LLM for GNN. LLMs extract additional semantic features or labels from textual data, serving as supervision signals for GNN training  (Chen et al., 2024c; Liu et al., 2024a; Zhu et al., 2024). However, current loosely coupled architectures with two-stage alignment struggle to fully exploit the synergy between GNNs and LLMs, resulting in suboptimal graph-text alignment in TAGs.
Recently, a noteworthy trend has emerged toward implementing LLM as GNN, where graph verbalizers convert graph data into code-like or heuristic prompts, enabling LLMs prompts to understand graph semantics and structure  (Ye et al., 2024; Wang et al., 2024a; Chen et al., 2024a). However, we argue that currently there are no true examples of this category, as they lack the essence of a genuine GNN: the message-passing paradigm. As shown in Figure 2(a), a traditional GNN layer includes neighbor sampling, aggregation-update, and optimization  (Kipf and Welling, 2017; Velickovic et al., 2017). By stacking multiple layers, structure-less embeddings are gradually transformed into structure-rich embeddings with higher-order signals. Due to the absence of these key components, an urgent challenge arises: Can we leverage LLMs to faithfully replicate GNNs to capture both graph semantics and structures simultaneously?
Meanwhile, existing works intuitively treat each graph node as an out-of-vocabulary (OOV) token, leading to graph-specific semantics and uncontrolled token explosion  (Tang et al., 2024; Ye et al., 2024). Worse still, due to vocabulary mismatches, ID-based node embeddings within graphs and language-based token embeddings from templates reside in different feature spaces, resulting in semantic misalignment in LLM inference. This incompatibility further hinders the transferability and scalability of graph-specific knowledge across other graphs and tasks. To enable knowledge transfer, a critical challenge emerges: Can we replace OOV tokens with compatible and universal node representations to build a versatile GFM?
A versatile GFM should be grounded in graph vocabulary learning  (Mao et al., 2024; Cai, 2024), ensuring the following properties: (1) Expressiveness: it encapsulates both semantic and structural information across diverse graphs. (2) Transferability: Every node in any graph should be representable using one or more fundamental units within the vocabulary. (3) Scalability: the vocabulary should be inclusive to accommodate unseen nodes, even those beyond existing graphs. Since natural language is a highly expressive medium made up of meaningful and transferable tokens  (Raffel et al., 2020; Radford et al., 2021; Palo et al., 2023; Wang et al., 2024c), we are inspired to establish a universal graph vocabulary within the text space for node representations. Thus, we propose PromptGFM as a GFM for TAGs as follows.
Figure 2: LLM-driven replication of the GNN workflow. We achieve fine-grained alignment between traditional embedding-based GNN and our prompt-based GNN.
Graph Understanding Module. To function LLMs as GNNs, we initialize node features with textual attributes and prompt LLMs to explicitly replicate the fine-grained GNN workflow within the text space. As illustrated in Figure 2, we represent graph structure through one-hop neighbor descriptions and design prompts to guide a flexible aggregation-update mechanism. To optimize, we incorporate heuristic prompts to reflect contrastive loss in each layer. Finally, iterative LLM calls simulate message passing, progressively refining verbose textual features into concise yet meaningful textual representations rather than numerical embeddings. Aligning with embedding-based GNN, our prompt-based GNN successfully preserves node semantics while capturing higher-order connections. Consequently, LLMs can operate as GNNs, and GNNs can be interpreted as LLMs, unlocking the full potential of GNN-LLM integration and empowering elegant graph-text alignment.
Graph Inference Module. Having captured semantic and structural information through prompt-based GNNs, we decouple textual representations to establish a graph vocabulary, where each node is mapped to a finite sequence of language-based tokens, essentially as language-based IDs. This vocabulary is universally transferable and scalable across graphs, which resolves semantic irrelevance and graph-text incompatibility. Afterward, these language-based IDs can be incorporated to generate readable and coherent pure-language prompts. Within a multi-instruction fine-tuning framework, we collect diverse instructions across graphs and tasks to effectively fine-tune an LLM, enabling cross-graph and cross-task knowledge transfer. In conclusion, this graph vocabulary eliminates incompatibility and paves the way for general GFMs.
2 Related Works
GNN-LLM Integration. LLMs have unlocked unprecedented potential for graph machine learning, driving efforts to integrate GNNs and LLMs for modeling TAGs. (a) GNN for LLM. As structure tokenizers, GNNs and graph transformers embed graph topology into textual representations, enhancing semantical understanding in LLMs (Tang et al., 2024; Chai et al., 2023; Liu et al., 2024b). However, effectively coordinating GNN-LLM architectures and co-training remains a challenge. (b) LLM for GNN. LLMs assist GNNs by generating node-level and edge-level labels to address data sparsity issues  (Chen et al., 2024c; Guo et al., 2024; Xia et al., 2024; Shu et al., 2024). Meanwhile, the other line directs LLMs to produce additional features or explanations to overcome semantic deficiencies  (Liu et al., 2024a; Zhu et al., 2024; He et al., 2024). However, their reliance on LLM-generated content inevitably introduces noise, impacting performance. (c) LLM as GNN. This paradigm directly operates LLMs as GNNs by designing structure verbalizers to convert graph data into code-like or heuristic prompts for LLM inference  (Chen et al., 2024a; Ye et al., 2024; Wang et al., 2024a, b). Yet, they fail to capture high-order connections due to the lack of an intrinsic GNN mechanism. Overall, the current decoupled approaches rely on two-stage alignment, failing to fully exploit the strengths of both models. This limitation motivates us to propose a new paradigm where LLMs inherently function as GNNs, maximizing their synergistic potential.
Graph-Text Alignment in Embedding Space. Current approaches to graph-text alignment primarily operate in the embedding space. One approach uses graph encoders as prefixes, mapping graph-aware embeddings to language-based embeddings for LLMs fine-tuning in a shared space  (Huang et al., 2023, 2024). Other works adopt a two-tower architecture, leveraging contrastive learning  (Li et al., 2023a; Brannon et al., 2023; Tang et al., 2024; Jin et al., 2024b), iterative training  (Zhao et al., 2023a; Zhu et al., 2024), or knowledge distillation  (Mavromatis et al., 2023) to align distinct representations. However, these methods encounter a persistent modality gap, limiting transferability and scalability across graphs. Since TAGs inherently contain textual information, we advocate for shifting graph-text alignment to the text space.
Graphs Foundation Models. A GFM aims to achieve transferability across different datasets and tasks, where the key challenge lies in finding a graph vocabulary that identifies transferable units to encode invariance on graphs  (Liu et al., 2023; Mao et al., 2024). Previous works rely on domain-specific vocabularies: GraphGPT  (Tang et al., 2024) assumes unique IDs to nodes and creates a dataset-specific vocabulary, while MoleBERT  (Xia et al., 2023) defines a molecular graph vocabulary by converting atomic properties into chemically meaningful codes. Despite their success, they lack in-context learning and cross-domain transferability. Recently, while some studies have investigated understanding and inferring graphs in natural language  (Fatemi et al., 2024; Wang et al., 2023; Zhao et al., 2023b; Liu et al., 2024a; Zhang et al., 2024b, a), none of them have attempted to establish a language-based graph vocabulary. To address this, we introduce an expressive graph vocabulary to exploit the inherent transferability of natural language, advancing the development of a versatile GFM.
3 Preliminaries
Text-Attributed Graphs. A TAG is formally represented as ${G}=\left(V,E,X\right)$, where $V$ is the set of nodes, $E$ is the set of edges. In this work, each node $v_{i}\in{V}$ is associated with a textual description ${X}_{i}=\left(x_{i}^{1},x_{i}^{2},\ldots,x_{i}^{{n_{i}}}\right)$, where each $x_{i}^{k}\in\mathcal{X},k=1,\dots,{n_{i}}$. Here, $X$ denotes the textual attributes for nodes, and $\mathcal{X}$ represents the natural language token dictionary.
Figure 3: The pipeline of PromptGFM. (a) Graph Understanding Module: For arbitrary TAGs from different domains, prompt-based GNN replicates traditional embedding-based GNN workflow in the text space, generating compact node representations. (b) Graph Inference Module: We establish a unified graph vocabulary and extract language-based IDs to generate massive pure-language prompts, enabling LLM fine-tuning across graphs and tasks.
Graph Neural Networks. GNNs have emerged as state-of-the-art models in graph machine learning, predominately relying on the message-passing paradigm. In practice, a GNN first selects neighboring nodes to a target node, aggregates their representations to capture local structure, and then updates the target node’s representation. Mathematically, for a given node $v_{i}$, the $l$-th layer of a general GNN can be formulated as:
$\mathcal{N}_{i}^{(l)}$ $=f_{\text{sample}}\left(\mathcal{N}_{i}\right),$ (1)
$\mathbf{m}_{i}^{(l)}$ $=f_{\text{agg}}\left({\left\{\mathbf{h}_{j}^{(l-1)},v_{j}\in\mathcal{N}_{i}^{(l)}\right\}}\right),$
$\mathbf{h}_{i}^{(l)}$ $=f_{\text{update}}\left(\mathbf{h}_{i}^{(l-1)},\mathbf{m}_{i}^{(l)}\right)$
where $\mathbf{h}_{i}^{(l)}$ is the node embedding of $v_{i}$ in the $l$-th layer. $\mathcal{N}_{i}$ denotes full neighbors and $\mathcal{N}_{i}^{(l)}$ is sampled neighbors in the $l$-th layer. To capture high-order connections, we stack $L$ layers and derive final embeddings as $\mathbf{h}_{i}=f_{\text{pooling}}\left(\mathbf{h}_{i}^{(1)},...,\mathbf{h}_{i}^{(L)}\right)$  (Grattarola et al., 2024). For optimization, a contrastive loss with negative sampling is commonly used in unsupervised graph learning  (Hamilton et al., 2017; Velickovic et al., 2019), expressed as:
$\begin{split}\ell=f_{\text{loss}}\left(f_{\text{sim}}\left(\mathbf{h}_{i},\mathbf{h}_{j}\right),f_{\text{sim}}\left(\mathbf{h}_{i},\mathbf{h}_{k}\right)\right),\end{split}$ (2)
where $v_{j}\in\mathcal{N}_{i}^{(l)}$ is a positive sample and $v_{k}\notin\mathcal{N}_{i}$ is a negative one. $f_{\text{sim}}\left(\cdot\right)$ measures the similarity between two nodes. $f_{\text{loss}}\left(\cdot\right)$ enforces contrastive learning by increasing similarity for connected nodes and reducing it for unconnected ones.
4 Methodology
This section describes the pipeline of our proposed PromptGFM, as illustrated in Figure 3.
4.1 Graph Understanding Module
The graph understanding module aims to generate expressive representations for each node within the graph, supporting the subsequent inference module. The main challenge lies in effectively aligning the semantic and structural information, where LLMs excel in textual understanding and GNNs in structural modeling  (Li et al., 2023b; Ren et al., 2024). To bridge the gap in GNN-LLM integration, we propose prompt-based GNNs that operate LLMs as GNNs by prompting LLMs to replicate general GNN workflow within the text space.
GNN Replication with LLMs. Our priority is to design appropriate prompts that enable LLMs to function as GNNs. This requires considering three essential factors: (a) Graph Representations: How can we effectively describe the node semantics and graph structure to LLMs? (b) Graph Structure: How can we incorporate message passing to encode structural dependencies? (c) Graph Semantics: How can we distill core semantics into concise yet expressive textual representations?
As illustrated in Figure 2, we perform a fine-grained replication of GNN, i.e. prompt-based GNN. First, we initialize node representations by summarizing textual attributes, akin to look-up layers in traditional GNNs. Similarly, we sample its one-hop neighbor information to reduce computational overhead as follows:
$\left\{X_{j}^{(l-1)},\left\{v_{j}\right\}\subset\mathcal{N}_{i}\right\}$ $\leftarrow\text{Prompt}_{\text{sample}}\left(X^{(l-1)},\mathcal{N}_{i}\right),$ (3)
where $X^{(l-1)}=\left\{X_{0}^{(l-1)},X_{1}^{(l-1)},\dots,X_{|V|-1}^{(l-1)}\right\}$ denotes textual representations of all nodes in the previous layer, and $\left\{X_{j}^{(l-1)}\right\}$ corresponds to the selected neighbors in $\mathcal{N}_{i}$. In practice, we limit neighbor sampling to a maximum of 20 to mitigate prompt length constraints. With this context, we directly prompt LLMs to perform essential message aggregation-update process, formulated as:
$X_{i}^{(l)}$ $\leftarrow\text{Prompt}_{\text{agg-upd}}\left(\left\{X_{j}^{(l-1)},\left\{v_{j}\right\}\in\mathcal{N}_{i}^{(l)}\right\},X_{i}^{(l-1)}\right),$ (4)
where we seek flexible message passing mechanism without specification, moving beyond traditional operators like mean and weighted aggregators. In terms of optimization, we adopt an unsupervised graph learning with contrastive loss to accommodate various downstream tasks, increasing neighbor similarity while separating distant nodes. Since negative sampling is redundant in the situation, we use LLM prompts to intuitively steer this process. Prompts can be found in Appendix D.
Figure 4: An instance of graph inference module in link prediction, where language-based IDs are indexed from the graph vocabulary to generate readable instructions using task-oriented templates. We adopt multi-instruction fine-tuning framework to unify diverse graphs and tasks and learn transferable knowledge for GFMs.
After repeating $L$ rounds, we obtain the final textual representation as $T_{i}=X_{i}^{(L)}$. These representations encapsulate both semantic and structural information, effectively solving the outlined issues: (a) Graph Representations: One-hop neighbor descriptions are equivalent to an adjacency matrix, representing the entire graph structure. (b) Graph Structure: To capture high-order relationships, we iteratively invoke LLMs using the same prompts, feeding each round’s output into the next. (c) Graph Semantics: Simultaneously, we explicitly instruct LLMs to produce concise yet expressive textual representations for each node, gradually refining them for denser and richer semantics.
We systematically examine how prompt-based GNNs faithfully mirror embedding-based GNNs in Appendix C. In summary, our fine-grained replication exemplifies the potential of using LLMs as GNNs, fostering seamless GNN-LLM integration and elegant graph-text alignment.
4.2 Graph Inference Module
The graph inference module seeks to unify diverse graphs and tasks and acquire transferable knowledge through multi-instruction fine-tuning using LLMs. Due to distinct vocabularies, existing methods, which represent nodes as OOV tokens, suffer from semantic misalignment between ID-based and language-based embeddings in task-oriented prompts, constraining their transferability. To tackle this limitation, we introduce a novel language-based graph vocabulary that bridges this incompatibility, enabling massive readable and coherent instructions to support LLM inference.
Graph Vocabulary Learning. Learning a transferable graph vocabulary, whose fundamental units can represent each node, is key to building GFMs. Its effectiveness depends on three essential criteria: expressiveness, transferability, and scalability. Since each node has been associated with a textual representation that encapsulates its core semantics and local structure, we intuitively propose a universal language-based graph vocabulary using these rich representations. In this vocabulary, each node is represented by a finite sequence of language tokens, i.e. a language-based ID, defined as follows:
$\mathcal{F}:V\rightarrow T,$ (5)
where each graph node $v_{i}$ is assigned a sequence as ${T}_{i}=\left(t_{i}^{1},t_{i}^{2},\ldots,t_{i}^{m_{i}}\right)$, where $t_{i}^{k}\in\mathcal{X}$ and $\mathcal{X}$ is a dictionary of general natural language tokens. Along this line, all nodes in $V$ have language-based IDs as $T=\left\{T_{0},T_{1},\dots,T_{|V|-1}\right\}$. Consequently, natural language tokens form the fundamental building blocks of our graph vocabulary, with structured token sequences representing graph nodes, analogous to words in standard lexicons.
Our graph vocabulary meets all expected criteria: (a) Expressiveness. Language-based IDs preserve rich semantic and structural information from an open-world setting. (b) Transferability. Like human lexicons, it shares natural language as a common foundation, ensuring inherent cross-graph transferability. (c) Scalability. Any node, whether previously seen or not, can dynamically generate its language-based ID, ensuring compatibility with existing nodes and mitigating token explosion.
Instruction Fine-Tuning. We employ a multi-instruction fine-tuning framework to incorporate various graphs and tasks (Chung et al., 2024; Wei et al., 2022). As illustrated in Figure 4, we index nodes from the graph vocabulary and embed their language-based IDs into task-oriented prompt templates to construct completed instructions $\mathcal{T}$:
$\mathcal{T}\leftarrow\text{Prompt}_{\text{template}}\left(T,G\mid\mathcal{F}\right),$ (6)
which include language-based IDs of the central nodes and their respective local structure $T$ from $G$, depending on specific graph-centric tasks (e.g., node classification or link prediction). Thus, these instructions are fully readable and coherent, as they are composed entirely of natural language tokens. Afterward, we convert all the question-answering tasks into a unified text-to-text format for LLM fine-tuning (Mishra et al., 2022). Given target sequences $Y$, the loss function is computed as:
$\mathcal{L}$ $=-\sum_{j=1}^{|Y|}\log\text{Pr}\left(Y_{j}\mid\mathcal{T},Y_{<j}\right),$ (7)
where $\text{Pr}\left(Y_{j}\mid\mathcal{T},Y_{<j}\right)$ is the probability of the $j$-th token $Y_{j}$ in the output sequence $Y$, conditioned on the instruction $\mathcal{T}$ and all previous tokens $Y_{<j}=(Y_{1},Y_{2},\dots,Y_{j-1})$. Using T5 (Raffel et al., 2020), FLAN  (Wei et al., 2022), or Llama  (Touvron et al., 2023) as the backbone, we can co-train a unified LLM across diverse graphs and tasks by fine-tuning, which enables the acquisition of open-world global knowledge and inclusive accommodation of unseen graphs or tasks.
Constrained Decoding via Prefix Tree Search. To mitigate LLM hallucination in link prediction, we introduce a constrained decoding method using a prefix tree search strategy to regulate LLM outputs  (Cao et al., 2021; Tan et al., 2024). Specifically, we craft a prefix tree from language-based IDs of all candidate nodes, where each tree node represents a natural language token. Each unique path from the root to a leaf corresponds to a specific language-based ID. During autoregressive generation, token output is restricted to a valid path within the prefix tree, ensuring predictions align with actual graph nodes and eliminating hallucinations. This effectiveness is attributed to the discrete nature of language-based IDs, further highlighting the flexibility of the proposed graph vocabulary.
[…]
Method Cora Citeseer PubMed Ogbn-arxiv History Photo WikiCS
MLP 62.29 64.42 62.88 62.07 64.62 61.21 68.41
GCN 82.47 76.11 77.36 66.15 81.93 78.58 76.33
GAT 82.92 77.30 74.36 65.29 82.85 82.38 78.21
SAGE 83.69 73.17 83.22 68.78 82.12 80.06 79.56
RevGNN 86.90 77.34 82.16 70.43 83.04 83.24 81.22
SGFormer 82.36 73.76 78.92 63.44 78.98 80.12 76.56
NodeFormer 81.55 72.98 76.49 73.21 79.60 78.51 75.47
BERT 79.02 72.83 76.74 71.90 72.97 68.82 77.98
Sentence-BERT 78.09 72.12 75.48 77.24 74.10 69.02 77.72
OFA 75.72 71.58 75.26 74.68 81.43 84.46 78.02
LLaGA 81.25 68.80 86.54 76.05 82.55 85.34 80.74
ENGINE 91.48 78.46 90.24 76.02 82.46 83.75 81.56
GraphPrompter 80.26 73.61 94.80 75.61 79.42 80.04 80.98
PromptGFM (Flan-T5) 91.72 84.49 92.83 80.58 82.33 85.41 81.49
PromptGFM (Llama3) 92.42 85.32 94.65 83.78 86.72 86.61 84.66
5 Experiments
[…]
Method Discriminative Setting
Cora Citeseer Obgn-arxiv PubMed
GCN 77.15 78.72 80.89 77.36
GAT 70.44 77.17 76.25 74.36
SAGE 85.31 87.15 80.76 83.22
GraphPrompter 90.10 91.67 73.21 80.49
Ours (Flan-T5) 90.57 92.03 81.12 87.64
Ours (Llama3) 91.68 93.46 84.27 88.12
In this section, we conduct extensive experiments to address the following research questions (RQs):
$\bullet$ RQ1: How does PromptGFM perform on supervised node classification and link prediction?
$\bullet$ RQ2: How is the transferability across diverse graphs and tasks as a versatile GFM?
$\bullet$ RQ3: How does each module contribute to overall performance of PromptGFM?
$\bullet$ RQ4: What affect the GNN replication?
Datasets. We present the following datasets covering three domains: (a) computer science: Cora (McCallum et al., 2000), Citeseer (Giles et al., 1998), Ogbn-arxiv (Hu et al., 2020), and WikiCS  (Mernyei and Cangea, 2020); (b) e-commerce: Photo and History (Ni et al., 2019; Yan et al., 2023); (c) biomedical: PubMed (Sen et al., 2008). Details can be found in Appendix A.
Baselines. We make comprehensive comparisons with existing methods in four categories: (1) Graph-agnostic methods. We consider the MLP model without graph structure. (2) GNN-based methods. We employ four fundamental GNN models: GCN  (Kipf and Welling, 2017), GAT  (Velickovic et al., 2017), GraphSAGE  (Hamilton et al., 2017), and ReVGNN  (Li et al., 2021). We also explore SGFormer  (Wu et al., 2023) and NodeFormer  (Wu et al., 2022), which leverage transformer architectures to model graph data. (3) BERT-based methods. We utilize BERT (Devlin et al., 2019) and Sentence-BERT  (Reimers and Gurevych, 2019) to obtain textual representations for downstream tasks. (4) GNN-LLM Integration methods. Following the proposed taxonomy, we select GraphPrompter  (Liu et al., 2024b) as an instance of using GNNs to enhance LLMs. OFA  (Liu et al., 2024a) and ENGINE  (Zhu et al., 2024) are examples of leveraging LLMs for GNNs. Besides, we incorporate LLaGA  (Chen et al., 2024a) as an attempt of implementing LLM as GNN. We provide baseline details in Appendix B.
Reproduction Settings. We implement PromptGFM in PyTorch and run all experiments on four NVIDIA RTX A6000 GPUs. The graph understanding module leverages OpenAI’s GPT-4o mini while we fine-tune a Flan-T5 or Llama3-8B in the graph inference module. For evaluation metrics, we use accuracy and Macro-F1 (abbreviated as M-F1) for node classification, and accuracy and HR@1 for link prediction, respectively. We provide further implementation details in Appendix E.
5.1 Performance Comparison (RQ1)
Main Results. We train each model independently from scratch on a single graph and compare their performance. Tables 1 and 2 present the results for node classification and link prediction accuracy, respectively. PromptGFM achieves substantial improvements over state-of-the-art models, leading to the following insights. First, graph-based models outperform graph-agnostic ones, highlighting the importance of structural information. Second, our method surpasses OFA, ENGINE, and GraphPrompter in node classification, partly due to their decoupled architectures in both GNN-for-LLM and LLM-for-GNN paradigms. Third, PromptGFM outperforms LLaGA, which relies on templates to understand graph semantics and structures. The results suggest that heuristic prompts alone fail to capture sufficient high-order signals without an actual GNN mechanism. In contrast, PromptGFM demonstrates the potential of leveraging LLMs as GNNs through a prompt-based GNN, establishing a novel paradigm for graph-text alignment in TAGs.
Generative Link Prediction. Following a transductive setting, we partition the graph by links, create an input graph from the training set, and finally predict unseen connections for test nodes, with existing nodes as potential candidates. From Table 3, PromptGFM consistently outperforms traditional GNN models. Unfortunately, existing GNN-LLM approaches overlook this setting due to their reliance on OOV token embeddings, preventing LLM outputs from mapping to specific nodes and leading to unresolved hallucination issues. In contrast, our framework represents nodes as finite token sequences, thereby allowing constrained decoding to regulate LLM outputs. This further underscores the critical role of our graph vocabulary and the flexibility and scalability of PromptGFM.
5.2 Cross-graph Transferability (RQ2)
We examine cross-graph transferability in both intra-domain and inter-domain scenarios. Following  (Chen et al., 2024b), we explore two settings: (1) Pre-training, where models transfer to unseen datasets without prior exposure during training, and (2) Co-training, where available target data is also integrated for joint training with auxiliary datasets. This results in four distinct transfer paradigms.
Intra-domain Cross-graph Transferability. As shown in Table 4, all settings surpass the off-the-shelf LLM. Meanwhile, incorporating Arxiv yields greater gains than Cora, suggesting that larger and richer datasets enhance transferability more effectively. In the pre-training setting, significant improvements over direct inference with an off-the-shelf LLM highlights strong zero-shot transferability within the computer science domain, while similar results on co-training further confirm effective knowledge transfer. These findings underscore the potential to collect more graph data and build more comprehensive and knowledgeable GFMs.
[…]
Method Generative Setting
Cora Citeseer PubMed
GCN 5.95 6.82 0.51
GAT 2.22 3.59 0.28
SAGE 6.59 8.73 0.45
PromptGFM 8.21 8.90 1.21
Inter-domain Cross-graph Transferability. We outline cross-domain results in Table 6. In the pre-training setting, while incorporating a single external graph enhances performance, adding multiple graphs can lead to negative transfer. More critically, in the co-training setting under supervised learning, performance deteriorates as more cross-domain data is introduced. This decline stems from domain-specific knowledge conflicts, potentially caused by catastrophic forgetting or incompatible hyperparameters that fail to accommodate to varying domain distributions. Although PromptGFM provides a unified interface for diverse graphs and tasks, we leave robust positive knowledge transfer as future work, further discussed in Section 7.
[…]
Table 4: Intra-domain cross-graph generalization. We transfer from the computer science domain to the Citeseer dataset for node classification task. The first row indicates direct inference using off-the-shelf LLM.
Source Pre-training Co-training
Cora Arxiv Acc($\uparrow$) M-F1($\uparrow$) Acc($\uparrow$) M-F1($\uparrow$)
27.64 17.10 84.49 82.31
51.63 45.10 84.96 83.22
60.34 54.81 85.45 83.91
61.25 55.66 86.77 84.24
[…]
Setting Cora Citeseer PubMed
Acc($\uparrow$) M-F1($\uparrow$) Acc($\uparrow$) M-F1($\uparrow$) Acc($\uparrow$) M-F1($\uparrow$)
zero-shot 18.54 12.16 27.64 17.10 39.12 39.84
[…]
ST-NC 91.72 90.06 84.49 80.13 90.67 91.82
5.3 Ablation Studies (RQ3)
To assess the contribution of each module, we introduce three ablated variants: (1) w/o understanding, which removes the prompt-based GNN; (2) w/o inference, which eliminates multi-instruction fine-tuning; and (3) w/o both, which directly prompts LLMs using raw textual attributes. Figure 5 reports node classification accuracy, where our full model consistently outperforms all variants. First, the drop in w/o understanding indicates that omitting GNN replication leads to a loss of semantic and structural information. Second, w/o inference shows a significant decline, emphasizing the crucial role of LLM fine-tuning in integrating domain-specific knowledge. Lastly, w/o both yields the worst results, underscoring the synergy between the understanding and inference modules.
Table 6: Inter-domain cross-graph generalization. We transfer from the computer science domain to biomedical domain (i.e. PubMed) on node classification.
Source Pre-training Co-training
Cora Citeseer Arxiv Acc($\uparrow$) M-F1($\uparrow$) Acc($\uparrow$) M-F1($\uparrow$)
39.12 39.84 90.67 91.82
51.76 52.84 86.34 87.29
40.12 42.38 85.21 86.12
60.21 62.02 82.17 86.01
50.17 51.74 84.39 81.94
57.28 59.71 81.32 83.14
55.34 57.11 82.13 83.10
53.07 54.90 80.43 80.79
5.4 Exploration Studies (RQ4).
We analyze factors affecting prompt-based GNNs. Since GNNs are permutation-invariant, we evaluate LLM-powered prompt-based GNNs using Flan-T5 with two variants: (1) shuffle nodes, which randomizes neighbor order, and (2) shuffle tokens, which injects position-specific tokens to nodes and shuffle them instead. As illustrated in Table 7, regardless of the shuffle method, node classification accuracy exhibits only minor fluctuations, having negligible impact on the stability of prompt-based GNNs.
Additionally, we investigate the impact of layer depth in Appendix F. Furthermore, we present a case study on layer-by-layer representations in Appendix G.1, and compare language-based IDs with keywords from citation datasets in Appendix G.2.
Figure 5: Ablation studies on node classification.
Table 7: Permutation sensitivity of prompt-based GNNs.
Variant Cora Citeseer
shuffle nodes 90.64 $\pm$1.12 83.79 $\pm$ 0.89
shuffle tokens 90.55 $\pm$1.03 84.28 $\pm$1.06
6 Conclusion
In this work, we present PromptGFM, a GFM for TAGs built on graph vocabulary learning. With GNN replication within the text space, we decouple refined textual node representations and establish a unified graph vocabulary. This vocabulary endows compatibility and scalability, facilitating effective LLM fine-tuning with readable instructions for enhanced transferability. Experiments validate its superior performance and effective cross-graph and cross-task generalization. Our research reveal the potential of using LLM as GNN, opening new avenues for developing GFMs in TAGs.
7 Limitations
The proposed PromptGFM framework introduces a promising paradigm for building a Graph Foundation Model (GFM) in Text-Attributed Graphs (TAGs). However, like any novel framework, it has limitations that should be acknowledged for future improvements. First, while we establish a foundation model interface, incorporating a large number of datasets may negatively impact transferability, particularly in cross-domain scenarios. A more stable knowledge transfer mechanism is needed to mitigate domain shifts and preserving generalization. Second, our prompt-based GNN processes each graph independently, disregarding dataset imbalances and cross-graph semantic similarities. One solution is to explore cross-graph semantic alignment and attentive learning for better knowledge integration across heterogeneous graph datasets. Third, our approach remains constrained to text-attributed graphs, leaving a gap in developing a fully text-free graph foundation model. These limitations highlight potential directions for future research in building more efficient and comprehensive GFMs.
APPENDIX
This appendix contains additional details for our paper, which is organized as follows:
  • §A provides Data Descriptions used in our experiments.
  • §B illustrates more details about Baselines employed for comparison.
  • §C shows systematic comparison between Embedding-based GNNs and Prompt-based GNNs.
  • §D analyzes Prompt Design in our PromptGFM framework.
  • §E shows more Implementation Details.
  • §F reports the Hyperparameter Sensitivity for prompt-based GNNs.
  • §G presents the Case Study to offer deeper insights of our research.
Appendix A Data Descriptions
Dataset Domain #Nodes #Edges #Classes Raw Text
Cora Computer Science 2708 10858 7 paper titles and abstracts
Citeseer Computer Science 3327 9464 6 paper titles and abstracts
Ogbn-arxiv Computer Science 169343 2332486 40 paper titles and abstracts
WikiCS Computer Science 11701 431726 10 wikipedia entry names and contents
Photo E-commerce 48362 873782 12 item titles and reviews
History E-commerce 41551 503180 12 item titles and descriptions
Pubmed Biomedical 19717 88648 3 paper titles and abstracts
Table 8: Statistics of seven benchmarking datasets from three domains.
We utilize seven public benchmarking datasets from three domains to evaluate our framework, and the statistics of these datasets are illustrated in Table 8. For consistency, all graphs are treated as undirected. Below, we provide detailed descriptions of each dataset.
A.1 Computer Science Datasets
$\bullet$ Cora (McCallum et al., 2000). The dataset includes a citation network consisting of 2,708 scientific publications in the field of machine learning, categorized into 7 classes based on their research topics. Nodes represent individual papers, and edges denote citation links between them, totaling 10,858 connections. Each paper is associated with its title and abstract.
$\bullet$ Citeseer (Giles et al., 1998). This dataset introduces a citation network comprising 3,327 scientific publications, categorized into 6 classes, including Agents, Artificial Intelligence, Database, Information Retrieval, Machine Learning, and Human-Computer Interaction. Nodes correspond to documents, and edges represent citation relationships between them, amounting to 9,464 links.
$\bullet$ obgn-arxiv (Hu et al., 2020). The ogbn-arxiv dataset is part of the Open Graph Benchmark and consists of a directed citation graph of 169,343 arXiv papers, categorized into 40 subject areas. Nodes represent individual papers, and edges indicate citation relationships, totaling over 2.3 million connections. Each paper includes textual data from its title and abstract.
$\bullet$ WikiCS (Mernyei and Cangea, 2020). The WikiCS dataset is a benchmark dataset derived from Wikipedia, designed for evaluating GNNs. It comprises a citation network with 11,701 nodes representing computer science articles and 431,726 edges corresponding to hyperlinks between them. The dataset features 10 classes corresponding to branches of computer science, with very high connectivity.
A.2 E-commerce Datasets
$\bullet$ History (Yan et al., 2023). The History dataset, extracted from the Amazon-Books dataset, includes items categorized under the second-level label "History". It comprises 41,511 nodes and 503,180 edges, where each node represents a book, and edges indicate frequent co-purchases or co-views between books. This dataset incorporates the title and description of each book as text attributes of the node. The classification task involves assigning books to 12 distinct categories.
$\bullet$ Photo (Yan et al., 2023). The Photo dataset, derived from the Amazon-Electronics dataset, consists of 48,362 nodes and 873,782 edges, forming a dense network that reflects user purchasing behavior. Each node represents an electronic product, while edges indicate frequent co-purchases or co-views between items. For textual attributes, user reviews are incorporated, prioritizing the most upvoted review when available; otherwise, a randomly selected review is used. The classification task involves categorizing electronic products into 12 distinct categories.
A.3 Biomedical Datasets
$\bullet$ PubMed (Sen et al., 2008). PubMed is a citation network of 19,717 scientific publications from the PubMed database pertaining to diabetes, classified into 3 classes: experimental induced diabetes, type 1 diabetes, and type 2 diabetes. Nodes are research papers, and edges signify citation links, amounting to 88,648 connections. This dataset is used for large-scale graph representation learning and evaluating algorithms in the biomedical domain.
Appendix B Baselines
We provide detailed information on the baseline models, categorized into: (1) Graph-agnostic methods, (2) GNN-based methods, (3) BERT-based methods, and (4) GNN-LLM integration methods.
B.1 Graph-agnostic methods.
$\bullet$ MLP. This method adopts a multi-layer perceptron to learn low-dimensional embeddings for each node. In our work, we randomly initialize node embeddings without textual attributes.
B.2 GNN-based methods.
$\bullet$ GCN  (Kipf and Welling, 2017). This model introduces a neural network architecture that generalizes convolution operations to graph-structured data, enabling effective semi-supervised learning by aggregating feature information from a node’s local neighborhood.
$\bullet$ GAT  (Velickovic et al., 2017). This method incorporates attention mechanisms into graph neural networks, allowing nodes to assign different importance weights to their neighbors during feature aggregation, which enhances performance by focusing on the most relevant connections.
$\bullet$ SAGE  (Hamilton et al., 2017). GraphSAGE is an inductive representation learning framework on large graphs; it generates node embeddings by sampling and aggregating features from a node’s local neighborhood, facilitating generalization to unseen nodes or graphs.
$\bullet$ ReVGNN  (Li et al., 2021). This method includes a recurrent graph neural network tailored for dynamic graphs, capturing temporal dependencies by updating node representations as events occur over time, which is crucial for modeling evolving graph structures.
$\bullet$ SGFormer  (Wu et al., 2023). This work introduces a transformer-based architecture designed for graph data, integrating spectral graph theory into the transformer framework. It aims to capture both local and global graph structures efficiently by incorporating spectral filters, enhancing the model’s ability to learn complex graph representations.
$\bullet$ NodeFormer  (Wu et al., 2022). This framework presents a scalable graph transformer model that utilizes a randomized attention mechanism to approximate full attention on graphs. By reducing computational complexity, it enables efficient learning on large-scale graphs while preserving the expressiveness of transformer architectures.
B.3 BERT-based methods.
$\bullet$ BERT  (Devlin et al., 2019). This work introduces a deep learning model that understands language context by processing text bidirectionally. It is pre-trained using masked language modeling (MLM) and next sentence prediction (NSP) to learn rich linguistic features. BERT can be fine-tuned for various NLP tasks, achieving state-of-the-art performance.
$\bullet$ Sentence-BERT  (Reimers and Gurevych, 2019). This model is a modification of BERT designed for sentence embeddings, allowing efficient comparison of semantic similarity. It uses a siamese or triplet network structure to generate fixed-size vector representations, making tasks like sentence similarity, clustering, and retrieval significantly faster. Unlike BERT, which requires computationally expensive cross-encoding, Sentence-BERT enables quick and effective comparisons using cosine similarity.
B.4 GNN-LLM integration methods.
$\bullet$ LLaGA  (Chen et al., 2024a). This model effectively integrates LLM capabilities to handle the complexities of graph-structured data. It transforms graph nodes into structure-aware sequences and maps them into token embedding space using a specialized projector. LLaGA excels in generalization and interpretability, performing strongly across various datasets and tasks. It also supports zero-shot learning, making it highly adaptable for unseen datasets.
$\bullet$ OFA  (Liu et al., 2024a). This paper proposes a framework that handles various graph classification tasks across different domains using a single model. It introduces the nodes-of-interest (NOI) subgraph mechanism to standardize different tasks with a single task representation. Additionally, a novel graph prompting paradigm to leverage in-context learning and apply the same architecture across diverse graph classification tasks, achieving generalization across multiple domains.
$\bullet$ GraphPrompter  (Liu et al., 2024b). This work introduces a novel framework designed to align graph with LLMs via soft prompts. Specifically, it adopts GNNs to capture graph structure and leverages an LLM to interpret the textual information at the node level. By prompt tuning, this approach demonstrates the potential of LLMs to effectively interpret graph structures, combining both semantic and structural insights for improved graph learning tasks.
$\bullet$ ENGINE  (Zhu et al., 2024). This paper proposes a parameter- and memory-efficient fine-tuning method for textual graphs by using LLMs as encoders. It combines the LLMs and GNNs through a tunable GNN-based side structure, called G-Ladder, alongside each LLM layer, effectively reducing training costs without compromising performance.
Figure 6: Graph understanding module via prompt-based GNNs. We prompt LLMs to achieve fine-grained reproduction of traditional GNN workflow, refining verbose textual representations into concise yet meaningful ones. In the prompt, neighbor sampling (see Equation 3) is highlighted in purple, the aggregation-update mechanism (see Equation 4) in blue, and the optimization in red.
Appendix C Embedding-based GNNs vs. Prompt-based GNNs
Table 6 illustrates the details of the graph understanding module, with the prompt-based GNN as its core component. To ensure clarity, we systematically compare traditional embedding-based GNNs with our proposed prompt-based GNN, emphasizing its advantages.
$\bullet$ Input and output. In the embedding-based GNN framework, for each node, structure-less embeddings are progressively refined to structure-rich embeddings, whereas verbose textual sequences are gradually converted to concise textual sequences in our prompt-based GNN.
$\bullet$ Message passing. The multi-layer embedding updates are mirrored by multi-round LLM calls in text space, both of which progressively refining the representations in different spaces.
$\bullet$ Neighbor sampling. The neighbor sampling operation used to reduce computational load in traditional embedding-based GNNs is analogous to the selected one-hop neighbor descriptions employed to address prompt length limitations in prompt-based GNNs.
$\bullet$ Aggregation-update mechanism. Embedding-based GNNs use predefined operators (e.g., mean aggregator, weighted aggregator, or LSTM aggregator) to achieve message passing in the embedding space, while prompt-based GNNs use straightforward prompts to guide LLMs in executing the process more flexibly without predefined rules.
$\bullet$ Optimization. In prompt-based GNNs, we use heuristic prompts at each layer to reflect the key idea of contrastive loss. These cumulative layer-by-layer prompts are comparable to the layer-wise loss combination, formally as mean pooling, commonly seen in embedding-based GNNs.
In conclusion, our approach enables LLMs to replicate the finest GNN workflow within the text space, fostering seamless graph-text alignment. By bridging structural and semantic information through language-based representations, this framework introduces a novel paradigm for GNN-LLM integration. It not only enhances the interpretability and transferability of graph learning but also unlocks new possibilities for leveraging LLMs in graph-related tasks without relying on numerical embeddings.
Appendix D Prompt Design
In this section, we provide the templates of prompts in our PromptGFM framework, including prompt-based GNNs and task-oriented prompts.
Prompt for node initialization. The title of the paper is <the title of the paper>, the abstract of the paper is <the abstract of the paper>. Please summarize the paper.
/*** This prompt varies depending on the dataset. The instance above is designed for citation datasets. ***/
Prompt for each layer of GNN replication. Given the central node <l-th round textual representation of the central node>. The selected one-hop neighbors are [< l-th round of node #1>, <l-th round node #2>, ... ,<l-th round node #N>]. Please aggregate neighbor nodes and update a concise yet meaningful representation for the central node. Note connected nodes should share similar semantics and vice versa.
Prompt for node classification. <the language-based ID of the central node> has 1-hop connections with [..., <language-based IDs of its 1-hop neighbors>, ...], and it also has 2-hop connections with [..., <language-based IDs of its 2-hop neighbors>, ...]. Which category should <the language-based ID of the central node> be classified as ?
Prompt for discriminative link prediction. <the language-based ID of the central node> has 1-hop connections with [..., <language-based IDs of its 1-hop neighbors>, ...], and it also has 2-hop connections with [..., <language-based IDs of its 2-hop neighbors>, ...]. Among <the language-based ID of the central node> and <the language-based ID of its negative sampling node>, which node will be connected to <the language-based ID of the central node> ?
Prompt for generative link prediction. <the language-based ID of the central node> has 1-hop connections with [..., <language-based IDs of its 1-hop neighbors>, ...], and it also has 2-hop connections with [..., <language-based IDs of its 2-hop neighbors>, ...]. Which node should be connected to <the language-based ID of the central node>?
Appendix E Implementation Details
We provide further information for reproduction. In the graph understanding module, we selected the number of layers for the prompt-based GNN from $\{1,2,3,4\}$. We randomly sampled 30% of the first-order neighbors during neighborhood sampling, capping the maximum number of sampled nodes at 20 to reduce computational cost and prevent overfitting. In the graph inference module, we fine-tuned the LLM with a learning rate of 3e-4 and a batch size of 4. To mitigate potential biases introduced by task-specific prompts, we designed a prompt pool for each task requirement and randomly selected prompts during instruction construction to enhance robustness. We employ 10-fold cross-validation and report average results across all folds. We employed a standard early-stopping strategy during training: if the performance metric on the validation set did not improve over a fixed number of consecutive epochs (determined based on the dataset), we halted training to prevent overfitting. For other hyperparameters of the compared methods, we referred to the original papers and carefully tuned them to suit each dataset. In terms of the TAG setting, we utilize the textual features for initialization in all embedding-based models.
Appendix F Hyperparameter Sensitivity
Figure 7: Impact of varying prompt-based GNN layers on node classification performance.
We explore the impact of the number of layers in our prompt-based GNN. As shown in Figure 7, we can observe that PromptGFM progressively improves as the layer increases due to its ability to capture broader context and higher-order relationships over the graph. However, after a certain point, further stacking layers results in diminishing returns or even performance degradation due to over-smoothing, where node representations become indistinguishable within their local structures. This trend is consistent with traditional GNNs. Optimal performance is achieved with 3-layer GNN for Cora, while Citeseer reaches its best results with 2 layers. This analysis suggests that textual representations can be propagated over the graph similarly to numerical embeddings, effectively capturing semantic and structural information simultaneously.
Appendix G Case Study
G.1 Textual Representations in Prompt-based GNNs
In this part, we select two representative cases in citation networks and demonstrate their layer-by-layer refinement of our prompt-based GNN. Specifically, we provide the textual representations at each layer, including Round 0 as initial features. More importantly, we also collect its one-hop nodes and highlighted the source and relevant information below. From our empirical studies, we have the following observations. Overall, it is evident that the verbose textual representations are progressively refined to concise textual presentations. Meanwhile, the core semantics become increasingly clear throughout the process, until a short sequence composed of several natural language tokens at the last round. Furthermore, we notice that we effectively incorporate the key ideas of some neighboring nodes, as reflected in the refined textual representations after each aggregation-update operation. More specifically, the neighboring information is naturally fused to the central node within the text space, boosting their connections and improving the expressiveness of the central node. This study further illustrates the explicit message passing through our proposed prompt-based GNN in the text space, which captures the graph semantics and structures simultaneously.
G.1.1 Case A
Raw Text Attributes. Paper Title: Evaluating Neural Network Predictors by Bootstrapping. Abstract: We present a new method, inspired by the bootstrap, whose goal it is to determine the quality and reliability of a neural network predictor. Our method leads to more robust forecasting along with a large amount of statistical information on forecast performance that we exploit.We exhibit the method in the context of multi-variate time series prediction on financial data from the New York Stock Exchange. It turns out that the variation due to different resamplings (i.e., splits between training, cross-validation, and test sets) is significantly larger than the variation due to different network conditions (such as architecture and initial weights). Furthermore, this method allows us to forecast a probability distribution, as opposed to the traditional case of just a single value at each time step. We demonstrate this on a strictly held-out test set that includes the 1987 stock market crash. We also compare the performance of the class of neural networks to identically bootstrapped linear models.
Round 0: Textual Feature Initialization. This method evaluates neural network predictors using a bootstrap-inspired approach, improving robustness in NYSE financial data forecasting. It predicts probability distributions rather than single values, surpassing resampling variations. Compared to bootstrapped linear models, it enhances accuracy, as demonstrated during the 1987 stock market crash.
Round 1: 1st Layer Representations. This bootstrap method improves neural network evaluations in NYSE forecasting, comparable to traditional error estimation methods. It clears and simplifies data structures, enhancing accuracy and predicting probability distribution.
/***"comparable to traditional error estimation methods" comes from the paper: A comparison of some error estimates for neural network models. ***/
/***"It clears and simplifies data structures" comes from the paper: The Observer-Observation Dilemma in Neuro-Forecasting: Reliable Models From Unreliable Data Through CLEARNING. ***/
Round 2: 2nd Layer Representations. This bootstrap-inspired method refines neural network evaluations for NYSE forecasting, highlighting local error bars from nonlinear regression analysis, improving accuracy with data simplification techniques.
/***"local error bars from nonlinear regression analysis" comes from the paper: Learning Local Error Bars for Nonlinear Regression. ***/
Round 3: 3rd Layer Representations. Bootstrap method refines neural forecasts, integrates error bars, enhances accuracy.
G.1.2 Case B
Raw Text Attributes. Paper Title: Insulin dependent diabetes in Asians. Abstract: Type 1 diabetes is said to be extremely rare in children in India, where diabetes treated with insulin may be due to chronic pancreatic disease or malnutrition. To see whether typical type 1 diabetes occurred in Asian children in the United Kingdom, all known Asian children with diabetes in industrial West Yorkshire were ascertained. A total of 17 such children were studied; of these, seven were from three multiplex families and two fathers from these families had diabetes. All children were ketosis prone and developed diabetes while resident in the UK. There were significant increases in HLA-B8 and HLA-DR3 and increases in HLA-DR4 and HLA-DR3/DR4, while HLA-B15 was absent. Islet cell antibodies, either IgG or complement fixing, were present in four of 18 subjects tested, all of whom had disease of short duration. The prevalence of type 1 diabetes in Asian children aged 15 years or less in West Yorkshire was 36/100,000, assuming complete ascertainment. It is concluded that typical type 1 diabetes may occur in Asian children and this condition may be more common in families who have migrated to the UK.
Round 0: Textual Feature Initialization. The study examined type 1 diabetes prevalence in 17 Asian children in West Yorkshire, UK, focusing on family history. Seven were from multiplex families, and all were ketosis-prone. Significant increases in HLA-B8, HLA-DR3, and HLA-DR4 were found, with no HLA-B15. Islet cell antibodies were present in four of 18 subjects. The diabetes prevalence was 36/100,000, suggesting it may be more common in migrant families.
Round 1: 1st Layer Representations. The study examined type 1 diabetes in Asian children in West Yorkshire, UK. Environmental factors and migration may raise incidence, unlike the low incidence of insulin-dependent diabetes in Karachi. Increases in HLA-B8, HLA-DR3, and HLA-DR4 were found.
/***"Environmental factors and migration may raise incidence" comes from the paper: Evidence for an environmental effect in the aetiology of insulin dependent diabetes in a transmigratory population. ***/
/***"low incidence of insulin-dependent diabetes in Karachi" comes from the paper: Incidence of insulin dependent diabetes mellitus in Karachi, Pakistan. ***/
Round 2: 2nd Layer Representations. The study examined type 1 diabetes in Asian children in West Yorkshire, UK. Migration may raise incidence, unlike the low incidence in Karachi. Increases in HLA markers were found, similar to North and South Indian diabetics.
/***"Migration may raise incidence" comes from the paper: HLA-DR antigen frequencies in a North Indian type I diabetic population. ***/
/***"North and South Indian diabetics" comes from two papers: HLA-DR antigen frequencies in a North Indian type I diabetic population and HLA, complement C2, C4, properdin factor B and glyoxalase types in South Indian diabetics. ***/
Round 3: 3rd Layer Representations. Type 1, diabetes. Migration and HLA markers linked to increased diabetes incidence.
G.2 Language-based IDs vs. Key Words
To provide further insights, we leverage external information to validate the superiority of PromptGFM in capturing the core semantics of nodes. Specifically, we extract key words of papers from citation datasets (Cora, Citeseer, Obgn-arxiv, and PubMed) and compare them with their language-based IDs in our universal graph vocabulary.
Table 9 summarizes the key words and language-based IDs of selected papers, along with their titles and URLs for reference. Overall, it is evident that there are strong semantic relevance between the language-based IDs and keywords. For example, regarding the paper titled Distributed Protocols at the Rescue for Trustworthy Online Voting, the key words have appeared within its language-based ID, suggesting that PromptGFM has effectively captured its core semantics through our prompt-based GNN. In addition, in Committees providing EJR can be computed efficiently, where the title is less indicative of the content, the language-based ID still aligns perfectly with the corresponding key words, such as efficient computation and rules. This finding demonstrates that PromptGFM not only effectively captures the core idea without relying on the title, but also filters relevant semantics from neighboring nodes to enhance its own representations. Overall, our language-based IDs accurately capture and extend the semantics of the nodes, making them well-suited to form a universal graph vocabulary.
Table 9: The comparison between language-based IDs generated by prompt-based GNNs as part of the graph vocabulary and keywords extracted from the original papers in citation datasets. To prevent data leakage, we exclude these keywords from model training. The observed similarity confirms that our prompt-based GNN effectively captures the essential semantics of nodes. Furthermore, our language-based IDs encode richer information than keywords by preserving high-order structural signals.
Paper Language-based ID Key Words URL
Modular Verification of Interrupt-Driven Software Modular verification of interrupt-driven software using abstract interpretation Software, Abstract Interpretation, Feasibility Verification arXiv:1709.10078
Parsimonious Data: How a single Facebook like predicts voting behaviour in multiparty systems Predicting voting behavior using Facebook likes in multiparty systems Facebook Likes, Voter Intention, Machine Learning, Multiparty System arXiv:1704.01143
A Fast Noniterative Algorithm for Compressive Sensing Using Binary Measurement Matrices Fast noniterative algorithm for compressive sensing with binary matrices Compressive Sensing, Deterministic Methods arXiv:1708.03608
Optimization of Battery Energy Storage to Improve Power System Oscillation Damping Battery storage optimization improves power system oscillation damping Battery Energy Storage System, Oscillation Damping arXiv:1811.10213
Neural Variational Hybrid Collaborative Filtering Neural Variational Hybrid Collaborative Filtering improves recommendation performance Collaborative Filtering, VAE, Recommendation System arXiv:1810.05376v6
Interpretable Neural Networks for Predicting Mortality Risk using Multi-modal Electronic Health Records Predicting mortality risk using interpretable multi-modal neural network Mortality Risk Prediction, Clinical Data arXiv:1901.08125
A New Approach to Distributed Hypothesis Testing and Non-Bayesian Learning: Improved Learning Rate and Byzantine-Resilience Distributed hypothesis testing with Byzantine resilience using Bayesian update Bayesian Learning, Byzantine Resilience arXiv:1907.03588
Accurate and Efficient Hyperbolic Tangent Activation Function on FPGA using the DCT Interpolation Filter Efficient hyperbolic tangent activation function using DCTIF Hyperbolic Tangent, Activation Function arXiv:1609.07750
Distributed Protocols at the Rescue for Trustworthy Online Voting Trustworthy online voting with distributed blockchain protocols Distributed Voting, Distributed Protocols arXiv:1705.04480
Committees providing EJR can be computed efficiently Efficient computation of approval-based multi-winner voting rules Approval-Based Voting, Multi-Winner Elections arXiv:1704.00356
Creatism: A deep-learning photographer capable of creating professional work Creatism: deep learning system for artistic photography creation Creatism, Evaluation of Photographic Quality, Deep Learning arXiv:1707.03491
Relation of familial patterns of coronary heart disease, stroke, and diabetes to subclinical atherosclerosis: the multi-ethnic study of atherosclerosis Family history beyond early-onset heart disease impacts atherosclerosis Family History, Coronary Heart Disease, Stroke doi.org/10.1097/GIM.0b0
13e31818e639b
Glycemic index, glycemic load, and risk of type 2 diabetes Benefits of low-GI diet in type 2 diabetes Diabetes, Prevention doi.org/10.1093/ajcn/76/1.
274S
Decreased insulin responsiveness of glucose uptake in cultured human skeletal muscle cells from insulin-resistant nondiabetic relatives of type 2 diabetic families Inherited defects contribute to insulin resistance in diabetes Insulin Resistance, Inherited Factors doi.org/10.2337/diabetes.
49.7.1169
Quantitative histopathological studies of the extramural coronary arteries from Type 2 (non-insulin-dependent) diabetic patients Histopathological study of coronary arteries in diabetic patients Histopathology, Diabetes Mellitus doi.org/10.1007/BF0027
4798
Metabolic control and diet in Finnish diabetic adolescents Factors influencing metabolic control in diabetic adolescents Diabetes Mellitus, Adolescent doi.org/10.1111/j.1651-2227.1992.tb12212.x