### v2

# LLM as GNN: Graph Vocabulary Learning for Text-Attributed Graph Foundation Models

Abstract
Text-Attributed Graphs (TAGs), where each node is associated with text descriptions, are ubiquitous in real-world scenarios. They typically exhibit distinctive structure and domain-specific knowledge, motivating the development of Graph Foundation Models (GFM) that generalize across diverse graphs and tasks. Despite large efforts to integrate Large Language Models (LLMs) and Graph Neural Networks (GNNs) for TAGs, most methods suffer from decoupled architectures with two-stage alignment, limiting their synergistic potential. Even worse, they assign out-of-vocabulary (OOV) tokens to graph nodes, leading to graph-specific semantics, token explosion, and incompatibility with task-oriented prompt templates, which hinders cross-graph and cross-task transferability. To address these challenges, we propose PromptGFM, a versatile GFM for TAGs grounded in graph vocabulary learning. PromptGFM comprises two key components: (a) Graph Understanding Module, which explicitly prompts LLMs to replicate the finest GNN workflow within the text space, facilitating seamless GNN-LLM integration and elegant graph-text alignment; (b) Graph Inference Module, which establishes a language-based graph vocabulary ensuring expressiveness, transferability, and scalability, enabling readable instructions for joint LLM fine-tuning. Extensive experiments demonstrate our superiority and transferability across diverse graphs and tasks. The code is available at https://github.com/agiresearch/PromptGFM.
1 Introduction
Graphs, characterized by their non-Euclidean structures and rich domain-specific knowledge, serve as fundamental representations of complex relational data. A wide range of real-world graphs further incorporate node-level textual information, forming Text-Attributed Graphs (TAGs), such as citation networks  [11, 18, 2], social networks  [22, 35], and molecular graphs  [52, 21]. However, existing solutions heavily rely on task- or dataset-specific training and deployment, hindering their transferability. To address this limitation, we seek to develop a Graph Foundation Model (GFM) capable of generalizing across different graphs and tasks for TAGs  [30, 56].
Figure 1: Overview of three GNN-LLM integration paradigms for graph-text alignment: (a) GNN for LLM and (b) LLM for GNN use decoupled architectures in the embedding space, while (c) our work functions LLM as GNN in the text space.
In TAGs, Graph Neural Networks (GNNs) and Large Language Models (LLMs) are employed to process different modalities, as illustrated in Figure 1. (a) GNN for LLM. GNNs generate structure-aware node embeddings that complement original textual embeddings, improving LLM inference  [44, 5, 29]. (b) LLM for GNN. LLMs extract additional semantic features or labels from textual data, serving as supervision signals for GNN training  [8, 27, 62]. However, existing two-stage alignment pipeline, where GNNs and LLMs are only loosely coupled, struggles to fully exploit their complementary strengths, often leading to suboptimal graph-text alignment.
Recently, a noteworthy trend has emerged toward using LLM as GNN, where graph verbalizers convert graph data into code-like or heuristic prompts, enabling LLMs to understand graph semantics and structure  [58, 49, 6]. However, we argue that these approaches do not qualify as true GNNs, as they omit the core message-passing mechanism. As illustrated in Figure 2(a), a traditional GNN layer includes neighbor sampling, aggregation-update, and optimization  [23, 46]. By stacking such layers, structure-less embeddings gradually evolve into structure-rich embeddings that encode higher-order dependencies. Without these essential components, a key challenge arises: Can we leverage LLMs to faithfully replicate GNNs to jointly capture both graph semantics and structure?
Meanwhile, existing works intuitively treat each graph node as an out-of-vocabulary (OOV) token, resulting in graph-specific semantics and uncontrolled vocabulary explosion  [44, 58]. Worse still, due to vocabulary mismatches, ID-based node embeddings and language-based token embeddings reside in different feature spaces, causing semantic misalignment during LLM inference. This incompatibility undermines both the transferability and scalability of graph-specific knowledge across graphs and tasks. To enable knowledge transfer, an urgent question emerges: Can we replace OOV tokens with compatible and universal node representations to build a versatile GFM?
Figure 2: LLM-driven replication of the GNN workflow. We achieve fine-grained alignment between traditional embedding-based GNN and our prompt-based GNN.
A versatile GFM should be grounded in graph vocabulary learning  [3, 30, 50] with the following desired properties: (1) Expressiveness. It encapsulates both semantic and structural information across diverse graphs. (2) Transferability. Every node in any graph should be representable by a composition of fundamental units within the vocabulary. (3) Scalability. The vocabulary should be inclusive to accommodate unseen nodes, even beyond existing graphs. Since natural language is inherently an expressive medium built upon meaningful and transferable tokens  [39, 38, 37, 50], we are inspired to establish a universal graph vocabulary in the text space for node representations. Therefore, we propose PromptGFM as a GFM for TAGs as follows.
Graph Understanding Module. To function LLMs as GNNs, we initialize node features with textual attributes and prompt LLMs to explicitly replicate the fine-grained GNN workflow in text space. As shown in Figure 2, we encode local structure via one-hop neighbor descriptions and design prompts that guide a flexible aggregation-update mechanism. To support optimization, we incorporate heuristic prompts that reflect contrastive loss at each layer. Finally, iterative LLM calls simulate the message passing over the whole graph, progressively refining verbose textual features into concise yet informative textual representations. Unlike traditional GNMs that output numerical embeddings, our prompt-based GNN produces structured text encodings that preserve both node semantics and high-order connections. Consequently, LLMs can operate as GNNs, and GNNs can be interpreted as LLMs, unlocking the full potential of GNN-LLM integration and empowering elegant graph-text alignment.
Graph Inference Module. Having captured semantic and structural information through prompt-based GNNs, we decouple the resulting textual representations to establish a graph vocabulary, where each node is mapped to a finite sequence of language-based tokens, essentially as language-based IDs. This vocabulary is universally interpretable, transferable across graphs, and scalable to unseen nodes, resolving semantic irrelevance and graph-text incompatibility. Afterward, language-based IDs can be incorporated to generate readable and coherent pure-language prompts. To support generalization, we adopt a multi-instruction fine-tuning framework that gathers diverse instructions from multiple graphs and tasks, facilitating cross-graph and cross-task knowledge transfer. In conclusion, this graph vocabulary eliminates incompatibility and paves the way for general GFMs.
2 Related Works
GNN-LLM Integration. LLMs have unlocked unprecedented potential for graph machine learning, driving efforts to integrate GNNs and LLMs for modeling TAGs. (a) GNN for LLM. As structure tokenizers, GNNs and graph transformers embed graph topology into textual representations, enhancing semantical understanding in LLMs [44, 5, 29]. However, effectively coordinating GNN-LLM architectures and co-training remains a challenge. (b) LLM for GNN. LLMs assist GNNs by generating node-level and edge-level labels to address data sparsity issues  [8, 15, 56]. Meanwhile, the other line directs LLMs to produce additional features or explanations to overcome semantic deficiencies  [27, 62, 17]. However, their reliance on LLM-generated content inevitably introduces noise, impacting performance. (c) LLM as GNN. This paradigm directly operates LLMs as GNNs by designing structure verbalizers to convert graph data into code-like or heuristic prompts for LLM inference  [6, 58, 49]. Yet, they fail to capture high-order connections due to the lack of an intrinsic GNN mechanism. Overall, the current decoupled approaches rely on two-stage alignment, failing to fully exploit the strengths of both models. This limitation motivates us to propose a new paradigm where LLMs inherently function as GNNs, maximizing their synergistic potential.
Graph-Text Alignment in Embedding Space. Current approaches to graph-text alignment primarily operate in the embedding space. One approach uses graph encoders as prefixes, mapping graph-aware embeddings to language-based embeddings for LLMs fine-tuning in a shared space  [19, 20]. Other works adopt a two-tower architecture, leveraging contrastive learning  [25, 1, 44], iterative training  [60, 62], or knowledge distillation  [31] to align distinct representations. However, these methods encounter a persistent modality gap, limiting transferability and scalability across graphs. Since TAGs contain textual information, we advocate for shifting graph-text alignment to the text space.
Graphs Foundation Models. A GFM aims to achieve transferability across different datasets and tasks, where the key challenge lies in finding a graph vocabulary that identifies transferable units to encode invariance on graphs  [28, 30]. Previous works rely on domain-specific vocabularies: GraphGPT  [44] assumes unique IDs to nodes and creates a dataset-specific vocabulary, while MoleBERT  [55] defines a molecular graph vocabulary by converting atomic properties into chemically meaningful codes. Despite their success, they lack in-context learning and cross-domain transferability. Recently, while some studies have investigated understanding and inferring graphs in natural language  [12, 48, 61, 27], none of them have attempted to establish a language-based graph vocabulary. To address this, we introduce an expressive graph vocabulary to exploit the inherent transferability of natural language, advancing the development of a versatile GFM.
3 Preliminaries
Text-Attributed Graphs (TAGs). A TAG is defined as ${G}=\left(V,E,X\right)$, where $V$ is the set of nodes, $E$ is the set of edges, and $X$ is the set of textual node attributes. Each node $v_{i}\in{V}$ is associated with a textual description ${X}_{i}=\left(x_{i}^{1},x_{i}^{2},\ldots,x_{i}^{{n_{i}}}\right)$, where each $x_{i}^{k}\in\mathcal{X},k=1,\dots,{n_{i}}$. Here, $\mathcal{X}$ denotes the natural language token dictionary.
Graph Neural Networks. GNNs have emerged as state-of-the-art models in graph learning, fundamentally built upon the message-passing mechanism. In practice, a GNN selects neighboring nodes to the target node, aggregates their representations to capture local structure, and updates the target node’s representation. Mathematically, for a node $v_{i}$, the $l$-th layer of a general GNN is defined as:
$\mathcal{N}_{i}^{(l)}$ $=f_{\text{sample}}\left(\mathcal{N}_{i}\right),$ (1)
$\mathbf{m}_{i}^{(l)}$ $=f_{\text{agg}}\left({\left\{\mathbf{h}_{j}^{(l-1)},v_{j}\in\mathcal{N}_{i}^{(l)}\right\}}\right),$
$\mathbf{h}_{i}^{(l)}$ $=f_{\text{update}}\left(\mathbf{h}_{i}^{(l-1)},\mathbf{m}_{i}^{(l)}\right)$
where $\mathbf{h}_{i}^{(l)}$ is the node embedding of $v_{i}$ at the $l$-th layer. $\mathcal{N}_{i}$ denotes full neighbors and $\mathcal{N}_{i}^{(l)}$ is sampled neighbors in the $l$-th layer. To capture high-order connections, we stack $L$ layers and derive final embeddings as $\mathbf{h}_{i}=f_{\text{pooling}}\left(\mathbf{h}_{i}^{(1)},...,\mathbf{h}_{i}^{(L)}\right)$  [14]. For optimization, a contrastive loss with negative sampling is commonly used in unsupervised graph learning  [16, 47], expressed as:
$\begin{split}\ell=f_{\text{loss}}\left(f_{\text{sim}}\left(\mathbf{h}_{i},\mathbf{h}_{j}\right),f_{\text{sim}}\left(\mathbf{h}_{i},\mathbf{h}_{k}\right)\right),\end{split}$ (2)
where $v_{j}\in\mathcal{N}_{i}^{(l)}$ is a positive sample and $v_{k}\notin\mathcal{N}_{i}$ is a negative one. $f_{\text{sim}}\left(\cdot\right)$ measures the similarity between two nodes. $f_{\text{loss}}\left(\cdot\right)$ enforces contrastive learning by increasing similarity for connected nodes and reducing it for unconnected ones.
4 Methodology
This section presents the pipeline of our proposed PromptGFM, which consists of a graph understanding module and a graph inference module, as illustrated in Figure 3.
4.1 Graph Understanding Module
The graph understanding module aims to generate expressive representations for each node within the graph to support downstream inference. The main challenge lies in effectively aligning semantic and structural information, where LLMs excel in textual understanding and GNNs in structural modeling  [26, 41]. To bridge the gap in GNN-LLM integration, we propose prompt-based GNNs that operate LLMs as GNNs by prompting LLMs to replicate general GNN workflow within the text space.
GNN Replication with LLMs. Our priority is to design appropriate prompts that enable LLMs to replicate GNNs. This requires considering three essential factors: (a) Graph Descriptions. How can we effectively describe the node semantics and graph structure to LLMs? (b) Graph Structure. How can we simulate message passing to encode structural dependencies? (c) Graph Semantics. How can we distill essential semantics into concise yet informative textual representations?
As illustrated in Figure 2, we implement a fine-grained replication of a general GNN workflow, i.e. prompt-based GNN. We begin by initializing node representations by summarizing their textual attributes, analogous to the lookup layer in conventional GNNs. To reduce computational overhead, we then sample one-hop neighbor information as follows:
$\left\{X_{j}^{(l-1)},\left\{v_{j}\right\}\subset\mathcal{N}_{i}\right\}$ $\leftarrow\text{Prompt}_{\text{sample}}\left(X^{(l-1)},\mathcal{N}_{i}\right),$ (3)
where $X^{(l-1)}=\left\{X_{0}^{(l-1)},X_{1}^{(l-1)},\dots,X_{|V|-1}^{(l-1)}\right\}$ denotes textual representations of all nodes from the previous layer, and $\left\{X_{j}^{(l-1)}\right\}$ corresponds to the selected neighbors $\mathcal{N}_{i}$. Then, we directly prompt LLMs to perform the core message aggregation-update process for each node, formulated as:
$X_{i}^{(l)}$ $\leftarrow\text{Prompt}_{\text{agg-update}}\left(\left\{X_{j}^{(l-1)},\left\{v_{j}\right\}\in\mathcal{N}_{i}^{(l)}\right\},X_{i}^{(l-1)}\right),$ (4)
where we seek flexible message passing mechanism without relying on predefined operators such as mean and weighted aggregators. A discussion on handling prompt length constraints is provided in Appendix F. In terms of optimization, we refer to unsupervised graph learning with contrastive loss to learn task-agnostic representations, which encourages similarity among neighboring nodes while pushing apart distant ones. Instead of traditional negative sampling, which becomes redundant in our setting, we leverage LLM prompts to intuitively steer the contrastive objective. Details of prompt construction are provided in Appendix D.
Figure 3: The pipeline of PromptGFM. (a) Graph Understanding Module: For arbitrary TAGs from different domains, prompt-based GNN replicates traditional embedding-based GNN workflow in the text space, generating compact node representations. (b) Graph Inference Module: We establish a unified graph vocabulary and extract language-based IDs to generate massive pure-language prompts, enabling LLM fine-tuning across graphs and tasks.
After repeating $L$ rounds, we obtain the final textual representation as $T_{i}=X_{i}^{(L)}$. These representations encapsulate both semantic and structural information, effectively solving the outlined issues: (a) Graph Description. One-hop neighbor descriptions are equivalent to an adjacency matrix, representing the entire graph structure. (b) Graph Structure. To capture high-order relationships, we iteratively invoke LLMs with the same prompts, feeding each round’s output into the next. (c) Graph Semantics. Simultaneously, we explicitly instruct LLMs to produce concise yet expressive textual representations for each node, gradually refining them for denser and richer semantics.
We systematically examine how prompt-based GNNs faithfully mirror embedding-based GNNs in Appendix C. In summary, our fine-grained replication exemplifies the potential of using LLMs as GNNs, fostering seamless GNN-LLM integration and elegant graph-text alignment.
4.2 Graph Inference Module
The graph inference module seeks to unify diverse graphs and tasks and acquire transferable knowledge through multi-instruction fine-tuning with LLMs. Existing methods, which represent nodes as OOV tokens, suffer from semantic misalignment between ID-based and language-based embeddings in task-oriented prompts due to distinct vocabularies, constraining their transferability. To tackle this limitation, we introduce a novel language-based graph vocabulary that bridges this incompatibility, enabling the generation of large-scale, coherent, and readable instructions to support LLM inference.
Graph Vocabulary. Learning a transferable graph vocabulary, whose fundamental units can represent each node, is critical for building effective GFMs. Its utility hinges on three key criteria: expressiveness, transferability, and scalability. Since each node has been equipped with a textual representation in the graph understanding module that captures its core semantics and local structure, we intuitively construct a universal language-based graph vocabulary based on these concise yet informative representations. In this vocabulary, each node is represented by a finite sequence of language tokens, known as a language-based ID, defined as:
$\mathcal{F}:V\rightarrow T,$ (5)
where each graph node $v_{i}$ is assigned a sequence as ${T}_{i}=\left(t_{i}^{1},t_{i}^{2},\ldots,t_{i}^{m_{i}}\right)$, where $t_{i}^{k}\in\mathcal{X}$ and $\mathcal{X}$ is a dictionary of general natural language tokens. Collectively, the language-based IDs for all nodes in $V$ are denoted as $T=\left\{T_{0},T_{1},\dots,T_{|V|-1}\right\}$. Along this line, natural language tokens serve as the fundamental building blocks of our graph vocabulary, with structured token sequences representing each graph node, analogous to words in standard lexicons.
Our vocabulary satisfies all expected criteria: (a) Expressiveness. Language-based IDs preserve rich semantic and structural information from open-world settings. (b) Transferability. Like human lexicons, the vocabulary is grounded in natural language, providing a shared foundation that ensures cross-graph interpretability and transferability. (c) Scalability. Any node, whether previously seen or not, can dynamically generate its language-based ID, maintaining compatibility with existing nodes while preventing token explosion. We analyze vocabulary coverage and uniqueness in Section 5.3.
Multi-Instruction Fine-Tuning. As shown in Figure 4, we employ a multi-instruction fine-tuning framework to incorporate diverse graphs and tasks [9, 51]. We index nodes from the vocabulary and embed their language-based IDs into task-oriented templates to construct completed instructions $\mathcal{T}$:
$\mathcal{T}\leftarrow\text{Prompt}_{\text{template}}\left(T,G\mid\mathcal{F}\right),$ (6)
which include language-based IDs of the central nodes, along with its corresponding local structure from $G$ depending on specific graph-centric tasks (e.g., node classification or link prediction). As these instructions are composed entirely of natural language tokens, they are fully readable and coherent. All graph-centric tasks are then reformulated in a text-to-text format, allowing the LLM to be fine-tuned within a unified instruction learning paradigm [34]. Given a target sequence $Y$, the loss function can be computed as:
$\mathcal{L}$ $=-\sum_{j=1}^{|Y|}\log\text{Pr}\left(Y_{j}\mid\mathcal{T},Y_{<j}\right),$ (7)
where $\text{Pr}\left(Y_{j}\mid\mathcal{T},Y_{<j}\right)$ denotes the probability of the $j$-th token $Y_{j}$ in the output sequence $Y$, conditioned on the instruction $\mathcal{T}$ and all previous tokens $Y_{<j}=(Y_{1},Y_{2},\dots,Y_{j-1})$. To support broad generalization, we fine-tune LLM backbones such as T5 [39], FLAN [51], and Llama [45] to co-train a unified model across diverse graphs and tasks. As a result, the model acquires open-world knowledge and generalizes effectively to previously unseen graphs and tasks.
Figure 4: An instance of graph inference module in link prediction, where language-based IDs are indexed from the graph vocabulary to generate readable instructions with task-oriented templates. A multi-instruction fine-tuning framework is employed to unify graphs and tasks and learn transferable knowledge for GFMs.
Constrained Decoding via Prefix Tree Search. To mitigate LLM hallucination, we introduce a constrained decoding strategy based on a prefix tree search to regulate LLM outputs  [4, 43]. Taking generative link prediction as an example, we craft a prefix tree from language-based IDs of all candidate nodes, where each tree node corresponds to a natural language token and each root-to-leaf path uniquely identifies a valid node ID. During autoregressive generation, token selection is constrained to a valid path within the prefix tree, ensuring that predictions align with actual graph nodes and eliminating hallucinations. This effectiveness can be attributed to the discrete nature of language-based IDs, further highlighting the flexibility of our proposed graph vocabulary.
5 Experiments
In this section, we conduct extensive experiments to address the following research questions (RQs):
$\bullet$ RQ1: How does PromptGFM perform on node classification and link prediction tasks?
$\bullet$ RQ2: How transferable is PromptGFM across diverse graphs and tasks as a general-purpose GFM?
$\bullet$ RQ3: How effective is the proposed graph vocabulary in terms of quantitative evaluation?
$\bullet$ RQ4: How does each module contribute to the overall performance of PromptGFM?
$\bullet$ RQ5: What is the computational efficiency in graph understanding and inference modules?
$\bullet$ RQ6: What factors influence the effectiveness of GNN replication within PromptGFM?
Datasets. We present the following datasets covering three domains: (a) computer science: Cora [32], Citeseer [13], Ogbn-arxiv [18], and WikiCS  [33]; (b) e-commerce: Photo and History [36, 57]; (c) biomedical: PubMed [42]. Details can be found in Appendix A.
Baselines. We make comprehensive comparisons with existing methods in four categories: (1) Graph-agnostic methods. We consider the MLP model without graph structure. (2) GNN-based methods. We employ four fundamental GNN models: GCN  [23], GAT [46], GraphSAGE [16], and ReVGNN [24]. We also explore SGFormer [54] and NodeFormer [53], which leverage Transformer architectures to model graph data. (3) LM-only methods. We utilize BERT [10], Sentence-BERT  [40], and Llama3-8B [45] to address downstream tasks. (4) GNN-LLM Integration methods. Following the proposed taxonomy, GraphPrompter  [29] is an instance of using GNNs to enhance LLMs. OFA  [27] and ENGINE  [62] are examples of leveraging LLMs for GNNs. Besides, we incorporate LLaGA  [6] as an attempt of implementing LLM as GNN. Details can be found in Appendix B.
[…]
Method Cora Citeseer PubMed Ogbn-arxiv History Photo WikiCS
MLP 62.29 64.42 62.88 62.07 64.62 61.21 68.41
GCN 82.47 76.11 77.36 66.15 81.93 78.58 76.33
GAT 82.92 77.30 74.36 65.29 82.85 82.38 78.21
SAGE 83.69 73.17 83.22 68.78 82.12 80.06 79.56
RevGNN 86.90 77.34 82.16 70.43 83.04 83.24 81.22
SGFormer 82.36 73.76 78.92 63.44 78.98 80.12 76.56
NodeFormer 81.55 72.98 76.49 73.21 79.60 78.51 75.47
BERT 79.02 72.83 76.74 71.90 72.97 68.82 77.98
Sentence-BERT 78.09 72.12 75.48 77.24 74.10 69.02 77.72
Llama3 78.17 74.68 77.46 77.65 76.49 74.42 79.94
OFA 75.72 71.58 75.26 74.68 81.43 84.46 78.02
LLaGA 81.25 68.80 86.54 76.05 82.55 85.34 80.74
ENGINE 91.48 78.46 90.24 76.02 82.46 83.75 81.56
GraphPrompter 80.26 73.61 94.80 75.61 79.42 80.04 80.98
PromptGFM (Flan-T5) 91.72 84.49 92.83 80.58 82.33 85.41 81.49
PromptGFM (Llama3-8B) 92.42 85.32 94.65 83.78 86.72 86.61 84.66
$p$-value 4e-3 2.73e-8 - 4.48e-10 4.85e-7 3.62e-2 9.93e-4
[…]
Method Discriminative Setting
Cora Citeseer Obgn-arxiv PubMed
GCN 77.15 78.72 80.89 77.36
GAT 70.44 77.17 76.25 74.36
SAGE 85.31 87.15 80.76 83.22
GraphPrompter 90.10 91.67 73.21 80.49
Ours (Flan-T5) 90.57 92.03 81.12 87.64
Ours (Llama3) 91.68 93.46 84.27 88.12
[…]
Method Generative Setting
Cora Citeseer PubMed
GCN 5.95 6.82 0.51
GAT 2.22 3.59 0.28
SAGE 6.59 8.73 0.45
PromptGFM 8.21 8.90 1.21
Reproduction Settings. We conduct all experiments in PyTorch on four NVIDIA RTX A6000 GPUs. The graph understanding module uses OpenAI’s GPT-4o mini while we fine-tune a Flan-T5 or Llama3-8B in the graph inference module. We employ 10-fold cross-validation and compute average results with standard deviations. For evaluation, we use accuracy and Macro-F1 (M-F1) for node classification, and accuracy and HR@1 for link prediction. To align with the TAG setting, BERT is used for node initialization across all applicable baselines. More details are provided in Appendix E.
5.1 Performance Comparison (RQ1)
Main Results. We train each model independently from scratch on a single graph and compare their performance. Tables 1 and 3 summarize the results for node classification and link prediction accuracy, respectively. PromptGFM achieves substantial improvements over state-of-the-art models, leading to the following observations. (a) Graph-based models consistently outperform graph-agnostic ones, highlighting the importance of structural information. (b) Our method surpasses OFA, ENGINE, and GraphPrompter on node classification, suggesting the limitations of their decoupled architectures in both GNN-for-LLM and LLM-for-GNN paradigms. (c) PromptGFM also surpasses LLaGA, which relies on templates to implicitly understand graph semantics and structures. These results indicate that heuristic prompting alone is insufficient to capture high-order dependencies without a true GNN mechanism. In contrast, PromptGFM effectively replicates message passing through a prompt-based GNN, demonstrating the feasibility of using LLMs as GNNs and establishing a new paradigm for graph-text alignment.
Generative Link Prediction. Following a transductive setting, we partition the graph by links, construct an input graph from the training edges, and predict missing connections for test nodes, using existing nodes as candidate targets. From Table 3, PromptGFM consistently outperforms traditional GNN models in the generative link prediction. Unfortunately, existing GNN-LLM methods fail to address this task, as their reliance on OOV token embeddings prevents LLM outputs from being directly mapped to specific nodes, often resulting in hallucinated predictions. In contrast, our work represents nodes as finite token sequences, allowing constrained decoding via prefix tree search to regulate LLM outputs. This further highlights the essence of our graph vocabulary and demonstrates the flexibility of PromptGFM in generative scenarios.
Table 4: Intra-domain cross-graph transferability
to Citeseer within the computer science domain.
Source Pre-training Co-training
Cora Arxiv Acc($\uparrow$) M-F1($\uparrow$) Acc($\uparrow$) M-F1($\uparrow$)
27.64 17.10 84.49 82.31
51.63 45.10 84.96 83.22
60.34 54.81 85.45 83.91
61.25 55.66 86.77 84.24
Table 5: Inter-domain cross-graph transferability from e-commerce to biomedical domain (i.e. PubMed).
Source Pre-training Co-training
History Photo Acc($\uparrow$) M-F1($\uparrow$) Acc($\uparrow$) M-F1($\uparrow$)
39.12 39.84 90.67 91.82
46.24 47.41 88.52 89.39
44.21 44.93 85.34 85.57
45.18 45.65 86.22 87.03
5.2 Cross-graph and Cross-task Transferability (RQ2)
We investigate cross-graph transferability in both intra-domain and inter-domain scenarios. Following  [7], we explore two settings: (1) Pre-training, where models are evaluated on entirely unseen datasets without prior exposure during training, and (2) Co-training, where a portion of the target dataset is available and jointly used with auxiliary datasets for training. These two dimensions yield four distinct transfer paradigms, enabling a comprehensive assessment of cross-graph transferability.
Intra-domain Cross-graph Transferability. As illustrated in Table 5, all settings surpass direct inference using an off-the-shelf LLM. Notably, incorporating Arxiv yields greater gains than Cora, suggesting that larger and richer datasets enhance transferability more effectively. In the pre-training setting, substantial improvements over direct inference demonstrate strong zero-shot transferability within the domain. A similar, albeit smaller, trend is observed in the co-training setting, further validating the effectiveness of intra-domain knowledge transfer. These findings underscore the potential to collect more graph data and build more knowledgeable and transferable GFMs.
Table 6: Inter-domain cross-graph transferability from computer science to biomedical domain (i.e. PubMed).
Source Pre-training Co-training
Cora Citeseer Arxiv Acc($\uparrow$) M-F1($\uparrow$) Acc($\uparrow$) M-F1($\uparrow$)
39.12 39.84 90.67 91.82
51.76 52.84 86.34 87.29
40.12 42.38 85.21 86.12
60.21 62.02 82.17 86.01
50.17 51.74 84.39 81.94
57.28 59.71 81.32 83.14
55.34 57.11 82.13 83.10
53.07 54.90 80.43 80.79
Inter-domain Cross-graph Transferability. We report cross-domain transfer results from e-commerce and computer science to the biomedical domain in Tables 5 and 6, respectively. In the pre-training setting, while incorporating a single external graph enhances performance, adding multiple source graphs may lead to negative transfer. More critically, in the co-training setting under supervised learning, performance consistently deteriorates as more cross-domain data is introduced. This decline stems from domain-specific knowledge conflicts, potentially caused by catastrophic forgetting or incompatible hyperparameters that fail to adapt to varying domain distributions [7, 59]. While PromptGFM offers a unified interface for diverse graphs and tasks, achieving robust and reliable cross-domain transfer remains a challenge, which we leave for future exploration.
[…]
Setting Cora Citeseer PubMed
Acc($\uparrow$) M-F1($\uparrow$) Acc($\uparrow$) M-F1($\uparrow$) Acc($\uparrow$) M-F1($\uparrow$)
zero-shot 18.54 12.16 27.64 17.10 39.12 39.84
[…]
ST-NC 91.72 90.06 84.49 80.13 90.67 91.82
[…]
5.3 Exploration Studies
Vocabulary Analysis (RQ3). We evaluate the quality of our graph vocabulary using two metrics in Table 8. (a) Coverage. We compute the proportion of nodes that can be matched to IDs. Since the vocabulary is language-based, any node’s semantics can, in principle, be expressed through a combination of language tokens. Thus, every node is successfully assigned a valid ID, confirming complete coverage. (b) Uniqueness. We measure the entropy of token distributions across nodes in both individual and merged graphs. Higher entropy indicates more diverse and fine-grained token usage, indicating stronger uniqueness and representational capacity. Notably, the merged vocabulary (“ALL”) achieves the highest entropy (12.2761), showing that integrating multiple graphs improves vocabulary expressiveness and discriminability, which reinforces the necessity of building GFMs.
Ablation Studies (RQ4). We construct three ablated variants to assess the contribution of each component: (a) w/o understanding, which removes the prompt-based GNN; (b) w/o inference, which eliminates multi-instruction fine-tuning; and (c) w/o both, which directly prompts LLMs using raw textual attributes. Figure 6 reports the node classification accuracy, where our full model consistently outperforms all variants. Specifically, the performance drop in w/o understanding highlights the importance of replicating GNN workflow to preserve both semantic and structural signals. The sharp decline in w/o inference further emphasizes the necessity of LLM fine-tuning for incorporating domain-specific knowledge. Finally, w/o both yields the worst results, underscoring the complementarity between the graph understanding and inference modules.
Figure 5: Ablation studies on node classification.
Figure 6: Impact of layer numbers.
Table 8: Quantitative graph vocabulary analysis on coverage and entropy.
Metric Cora Citeseer PubMed Ogbn-arxiv History Photo WikiCS ALL
Coverage (%) 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00
Entropy 9.1116 9.4561 8.4643 10.7464 10.3752 11.3272 9.3825 12.2761
Table 9: Efficiency analysis in terms of API cost (USD), number of tokens (M), training time (min), and power consumption (kWh).
Source Understanding Inference
Cora Citeseer Photo API Cost # Tokens Time Power
0.188 1.25 16.47 0.07
0.189 1.29 21.41 0.08
1.213 8.08 83.67 0.34
N/A N/A 35.31 0.16
N/A N/A 117.56 1.04
Efficiency Analysis (RQ5). We evaluate the computational and monetary cost of PromptGFM. As summarized in Table 9, while PromptGFM incurs an initial API cost during the graph understanding phase, this overhead is one-time and remains manageable with modern infrastructure. Specifically, each graph is processed only once to generate language-based node representations, which can be cached and reused throughout the inference pipeline. As a result, when additional graphs are introduced for training, there is no need to repeatedly call the LLM API to regenerate node representations. In contrast, traditional end-to-end models often require retraining from scratch when the graph structure changes, leading to higher cumulative cost. These characteristics make the computational overhead acceptable and amortizable, especially in dynamic or large-scale graph scenarios, where repeated retraining would otherwise be prohibitively expensive.
Prompt-based GNN Analysis (RQ6). We examine how the number of layers in our prompt-based GNN affects performance. As shown in Figure 6, increasing the number of layers initially improves results by capturing broader context and higher-order dependencies. However, beyond a certain depth, performance plateaus or declines due to over-smoothing, a trend consistent with traditional GNNs. Optimal performance is achieved with 3 layers on Cora and 2 layers on Citeseer. This suggests that textual representations can be propagated over the graph similarly to numerical embeddings, effectively capturing both semantics and structure. We further provide a case study on layer-wise representations in Appendix H.1, and compare language-based IDs with citation keywords in Appendix H.2. Additionally, we investigate permutation sensitivity of the prompts in Appendix G.
6 Conclusion
In this work, we present PromptGFM, a GFM for TAGs built on graph vocabulary learning. With GNN replication within the text space, we decouple refined textual node representations and establish a unified graph vocabulary. This vocabulary endows compatibility and scalability, facilitating effective LLM fine-tuning with readable instructions for enhanced transferability. Experiments validate its superior performance and effective cross-graph and cross-task generalization. Our research reveal the potential of using LLM as GNN, opening new avenues for developing GFMs in TAGs.
APPENDIX
This appendix contains additional details for our paper, which is organized as follows:
$\bullet$ §A provides Data Descriptions used in our experiments.
$\bullet$ §B illustrates more details about Baselines employed for comparison.
$\bullet$ §C shows comparison between Embedding-based and Prompt-based GNNs.
$\bullet$ §D analyzes Prompt Design in our PromptGFM framework.
$\bullet$ §E shows more Implementation Details.
$\bullet$ §F provides a Prompt Length Limitation Discussion.
$\bullet$ §G investigates Permutation Sensitivity in prompt-based GNNs.
$\bullet$ §H presents the Case Study to offer deeper insights of our research.
Appendix A Data Descriptions
Dataset Domain #Nodes #Edges #Classes Raw Text
Cora Computer Science 2708 10858 7 paper titles and abstracts
Citeseer Computer Science 3327 9464 6 paper titles and abstracts
Ogbn-arxiv Computer Science 169343 2332486 40 paper titles and abstracts
WikiCS Computer Science 11701 431726 10 wikipedia entry names and contents
Photo E-commerce 48362 873782 12 item titles and reviews
History E-commerce 41551 503180 12 item titles and descriptions
Pubmed Biomedical 19717 88648 3 paper titles and abstracts
Table 10: Statistics of seven benchmarking datasets from three domains.
We utilize seven public benchmarking datasets from three domains to evaluate our framework, and the statistics of these datasets are illustrated in Table 10. For consistency, all graphs are treated as undirected. Below, we provide detailed descriptions of each dataset.
A.1 Computer Science Datasets
$\bullet$ Cora [32]. The dataset includes a citation network consisting of 2,708 scientific publications in the field of machine learning, categorized into 7 classes based on their research topics. Nodes represent individual papers, and edges denote citation links between them, totaling 10,858 connections. Each paper is associated with its title and abstract.
$\bullet$ Citeseer [13]. This dataset introduces a citation network comprising 3,327 scientific publications, categorized into 6 classes, including Agents, Artificial Intelligence, Database, Information Retrieval, Machine Learning, and Human-Computer Interaction. Nodes correspond to documents, and edges represent citation relationships between them, amounting to 9,464 links.
$\bullet$ obgn-arxiv [18]. The ogbn-arxiv dataset is part of the Open Graph Benchmark and consists of a directed citation graph of 169,343 arXiv papers, categorized into 40 subject areas. Nodes represent individual papers, and edges indicate citation relationships, totaling over 2.3 million connections. Each paper includes textual data from its title and abstract.
$\bullet$ WikiCS [33]. The WikiCS dataset is a benchmark dataset derived from Wikipedia, designed for evaluating GNNs. It comprises a citation network with 11,701 nodes representing computer science articles and 431,726 edges corresponding to hyperlinks between them. The dataset features 10 classes corresponding to branches of computer science, with very high connectivity.
A.2 E-commerce Datasets
$\bullet$ History [57]. The History dataset, extracted from the Amazon-Books dataset, includes items categorized under the second-level label "History". It comprises 41,511 nodes and 503,180 edges, where each node represents a book, and edges indicate frequent co-purchases or co-views between books. This dataset incorporates the title and description of each book as text attributes of the node. The classification task involves assigning books to 12 distinct categories.
$\bullet$ Photo [57]. The Photo dataset, derived from the Amazon-Electronics dataset, consists of 48,362 nodes and 873,782 edges, forming a dense network that reflects user purchasing behavior. Each node represents an electronic product, while edges indicate frequent co-purchases or co-views between items. For textual attributes, user reviews are incorporated, prioritizing the most upvoted review when available; otherwise, a randomly selected review is used. The classification task involves categorizing electronic products into 12 distinct categories.
A.3 Biomedical Datasets
$\bullet$ PubMed [42]. PubMed is a citation network of 19,717 scientific publications from the PubMed database pertaining to diabetes, classified into 3 classes: experimental induced diabetes, type 1 diabetes, and type 2 diabetes. Nodes are research papers, and edges signify citation links, amounting to 88,648 connections. This dataset is used for large-scale graph representation learning and evaluating algorithms in the biomedical domain.
Appendix B Baselines
We provide detailed information on the baseline models, categorized into: (1) Graph-agnostic methods, (2) GNN-based methods, (3) LM-only methods, and (4) GNN-LLM integration methods.
B.1 Graph-agnostic methods.
$\bullet$ MLP. This method adopts a multi-layer perceptron to learn low-dimensional embeddings for each node. In our work, we randomly initialize node embeddings without textual attributes.
B.2 GNN-based methods.
$\bullet$ GCN  [23]. This model introduces a neural network architecture that generalizes convolution operations to graph-structured data, enabling effective semi-supervised learning by aggregating feature information from a node’s local neighborhood.
$\bullet$ GAT  [46]. This method incorporates attention mechanisms into graph neural networks, allowing nodes to assign different importance weights to their neighbors during feature aggregation, which enhances performance by focusing on the most relevant connections.
$\bullet$ SAGE  [16]. GraphSAGE is an inductive representation learning framework on large graphs; it generates node embeddings by sampling and aggregating features from a node’s local neighborhood, facilitating generalization to unseen nodes or graphs.
$\bullet$ ReVGNN  [24]. This method includes a recurrent graph neural network tailored for dynamic graphs, capturing temporal dependencies by updating node representations as events occur over time, which is crucial for modeling evolving graph structures.
$\bullet$ SGFormer  [54]. This work introduces a transformer-based architecture designed for graph data, integrating spectral graph theory into the transformer framework. It aims to capture both local and global graph structures efficiently by incorporating spectral filters, enhancing the model’s ability to learn complex graph representations.
$\bullet$ NodeFormer  [53]. This framework presents a scalable graph transformer model that utilizes a randomized attention mechanism to approximate full attention on graphs. By reducing computational complexity, it enables efficient learning on large-scale graphs while preserving the expressiveness of transformer architectures.
B.3 LM-only methods.
$\bullet$ BERT  [10]. This work introduces a deep learning model that understands language context by processing text bidirectionally. It is pre-trained using masked language modeling (MLM) and next sentence prediction (NSP) to learn rich linguistic features. BERT can be fine-tuned for various NLP tasks, achieving state-of-the-art performance.
$\bullet$ Sentence-BERT  [40]. This model is a modification of BERT designed for sentence embeddings, allowing efficient comparison of semantic similarity. It uses a siamese or triplet network structure to generate fixed-size vector representations, making tasks like sentence similarity, clustering, and retrieval significantly faster. Unlike BERT, which requires computationally expensive cross-encoding, Sentence-BERT enables quick and effective comparisons using cosine similarity.
$\bullet$ LLaMA3  [27]. LLaMA3-8B is an open-weight transformer model with 8 billion parameters, part of Meta’s third-generation LLaMA series. It is trained on 15T tokens from a diverse mix of web, code, and academic data, using a decoder-only architecture with a context window of 8,192 tokens. LLaMA3-8B achieves strong performance on a wide range of benchmarks, particularly in reasoning, multilingual tasks, and instruction following. Compared to earlier versions, it benefits from improved training stability, data quality, and scaling efficiency, making it a competitive foundation model for downstream adaptation and alignment.
B.4 GNN-LLM integration methods.
$\bullet$ LLaGA  [6]. This model effectively integrates LLM capabilities to handle the complexities of graph-structured data. It transforms graph nodes into structure-aware sequences and maps them into token embedding space using a specialized projector. LLaGA excels in generalization and interpretability, performing strongly across various datasets and tasks. It also supports zero-shot learning, making it highly adaptable for unseen datasets.
$\bullet$ OFA  [27]. This paper proposes a framework that handles various graph classification tasks across different domains using a single model. It introduces the nodes-of-interest (NOI) subgraph mechanism to standardize different tasks with a single task representation. Additionally, a novel graph prompting paradigm to leverage in-context learning and apply the same architecture across diverse graph classification tasks, achieving generalization across multiple domains.
$\bullet$ GraphPrompter  [29]. This work introduces a novel framework designed to align graph with LLMs via soft prompts. Specifically, it adopts GNNs to capture graph structure and leverages an LLM to interpret the textual information at the node level. By prompt tuning, this approach demonstrates the potential of LLMs to effectively interpret graph structures, combining both semantic and structural insights for improved graph learning tasks.
$\bullet$ ENGINE  [62]. This paper proposes a parameter- and memory-efficient fine-tuning method for textual graphs by using LLMs as encoders. It combines the LLMs and GNNs through a tunable GNN-based side structure, called G-Ladder, alongside each LLM layer, effectively reducing training costs without compromising performance.
Figure 7: Graph understanding module via prompt-based GNNs. We prompt LLMs to achieve fine-grained reproduction of traditional GNN workflow, refining verbose textual representations into concise yet meaningful ones. In the prompt, neighbor sampling (see Equation 3) is highlighted in darkred, the aggregation-update mechanism (see Equation 4) in blue, and the optimization in red.
Appendix C Embedding-based GNNs vs. Prompt-based GNNs
Figure 7 illustrates the details of the graph understanding module, with the prompt-based GNN as its core component. To ensure clarity, we systematically compare traditional embedding-based GNNs with our proposed prompt-based GNN, emphasizing its advantages.
$\bullet$ Input and output. In the embedding-based GNN framework, for each node, structure-less embeddings are progressively refined to structure-rich embeddings, whereas verbose textual sequences are gradually converted to concise textual sequences in our prompt-based GNN.
$\bullet$ Message passing. The multi-layer embedding updates are mirrored by multi-round LLM calls in text space, both of which progressively refining the representations in different spaces.
$\bullet$ Neighbor sampling. The neighbor sampling operation used to reduce computational load in traditional embedding-based GNNs is analogous to the selected one-hop neighbor descriptions employed to address prompt length limitations in prompt-based GNNs.
$\bullet$ Aggregation-update mechanism. Embedding-based GNNs use predefined operators (e.g., mean aggregator, weighted aggregator, or LSTM aggregator) to achieve message passing in the embedding space, while prompt-based GNNs use straightforward prompts to guide LLMs in executing the process more flexibly without predefined rules.
$\bullet$ Optimization. In prompt-based GNNs, we use heuristic prompts at each layer to reflect the key idea of contrastive loss. These cumulative layer-by-layer prompts are comparable to the layer-wise loss combination, formally as mean pooling, commonly seen in embedding-based GNNs.
In conclusion, our approach enables LLMs to replicate the finest GNN workflow within the text space, fostering seamless graph-text alignment. By bridging structural and semantic information through language-based representations, this framework introduces a novel paradigm for GNN-LLM integration. It not only enhances the interpretability and transferability of graph learning but also unlocks new possibilities for leveraging LLMs in graph-related tasks without relying on numerical embeddings.
Appendix D Prompt Design
In this section, we provide the templates of prompts in our PromptGFM framework, including prompt-based GNNs and task-oriented prompts.
Prompt for node initialization. The title of the paper is <the title of the paper>, the abstract of the paper is <the abstract of the paper>. Please summarize the paper.
/*** This prompt varies depending on the dataset. The instance above is designed for citation datasets. ***/
Prompt for each layer of GNN replication. Given the central node <l-th round textual representation of the central node>. The selected one-hop neighbors are [< l-th round of node #1>, <l-th round node #2>, ... ,<l-th round node #N>]. Please aggregate neighbor nodes and update a concise yet meaningful representation for the central node. Note connected nodes should share similar semantics and vice versa.
Prompt for node classification. <the language-based ID of the central node> has 1-hop connections with [..., <language-based IDs of its 1-hop neighbors>, ...], and it also has 2-hop connections with [..., <language-based IDs of its 2-hop neighbors>, ...]. Which category should <the language-based ID of the central node> be classified as ?
Prompt for discriminative link prediction. <the language-based ID of the central node> has 1-hop connections with [..., <language-based IDs of its 1-hop neighbors>, ...], and it also has 2-hop connections with [..., <language-based IDs of its 2-hop neighbors>, ...]. Among <the language-based ID of the central node> and <the language-based ID of its negative sampling node>, which node will be connected to <the language-based ID of the central node> ?
Prompt for generative link prediction. <the language-based ID of the central node> has 1-hop connections with [..., <language-based IDs of its 1-hop neighbors>, ...], and it also has 2-hop connections with [..., <language-based IDs of its 2-hop neighbors>, ...]. Which node should be connected to <the language-based ID of the central node>?
Appendix E Implementation Details
We provide additional details to facilitate reproducibility. In the graph understanding module, the number of layers in the prompt-based GNN is selected from ${1,2,3,4}$, and the first-order neighbor sampling ratio from ${0.3,0.6,1.0}$. In the graph inference module, we fine-tune the LLM with a learning rate of 3e-4 and a batch size of 4. To mitigate potential biases introduced by task-specific prompts, we designed a prompt pool for each task requirement and randomly selected prompts during instruction construction to enhance robustness. We employed a standard early-stopping strategy during training: if the performance metric on the validation set did not improve over a fixed number of consecutive epochs (determined based on the dataset), we halted training to prevent overfitting. For other hyperparameters of the compared methods, we referred to the original papers and carefully tuned them to suit each dataset.
Appendix F Prompt Length Limitation Discussion
In dense TAGs, high-degree nodes pose a challenge for prompt-based GNNs due to potentially excessive prompt lengths, which can exceed the input limits of current LLMs and impair model efficiency. To address this issue, we analyze in the following aspects that balances efficiency, scalability, and representation quality.
F.1 Neighbor Sampling in Scalable GNNs
Following standard GNN practices, we sample a fixed number of neighbors per node to reduce prompt length and improve robustness. As shown in Table 11, varying the sampling ratio across three datasets reveals that selecting 60% of neighbors yields the best performance, confirming its effectiveness. Meanwhile, limiting the maximum number of sampled neighbors is a common practice in scalable GNNs. For example, GraphSAGE samples 25 and 10 neighbors at the first and second hops, respectively, while PinSAGE selects from $\{10,20,50\}$. Inspired by these practices, we choose from $\{10,20,30,40\}$ and empirically set 20 as the cap, ensuring prompts remain within LLM input limits while preserving sufficient structural context.
F.2 Token Capacity of LLMs
Our implementation leverages GPT-4o mini, which supports up to 128k input tokens per call. This capacity allows a single prompt to encode information from over 300 neighbors, given that each neighbor’s textual summaries occupies approximately 400 tokens on average (for the first round of message propagation). As shown in Table 12, only a small fraction of nodes (<0.25%) in a few datasets exceed this limit. For these rare cases, we have applied neighbor sampling and truncation in Section F.1, effectively bounding prompt length without degrading representation quality.
F.3 Prompt Design
In our prompt-based GNN framework, we design prompts that guide the LLM to aggregate information from a node’s sampled neighbors and produce a concise yet expressive representation. This process is inherently iterative: with each layer, the prompt becomes more compressed as semantic and structural information is distilled into fewer tokens. Empirically, the average prompt length decreases from approximately 400 tokens in the first layer, to  125 tokens in the second, and eventually to  10 tokens per node on the Cora dataset.
Importantly, each node is processed independently, and its prompt length depends solely on its local neighborhood, i.e., its degree, rather than on global graph size. This localized design ensures scalability, as prompt lengths remain bounded even in large graphs, allowing our approach to maintain efficiency without sacrificing representational capacity.
Table 11: The performance with varying neighbor sampling ratios acorss three datasets.
Sampling Ratio 100% 60% 30%
Cora 81.36 89.35 88.36
Citeseer 79.34 82.23 81.13
PubMed 90.23 91.12 89.34
Table 12: Analysis of High-Degree Nodes and Average Connectivity
Metric Cora Citeseer PubMed Ogbn-arxiv History Photo WikiCS
Avg. Degree 4.01 2.84 13.77 36.90 18.07 12.11 4.50
High-degree (>300) Nodes 0 0 0 0.20% 0.25% 0.19% 0
Appendix G Permutation Sensitivity
We evaluate the sensitivity of our prompt-based GNNs (using Flan-T5) to input order by evaluating two variants: (a) shuffle nodes, which randomizes the order of neighbors, and (b) shuffle tokens, which injects position-specific tokens to nodes and shuffle them instead. As illustrated in Table 13, regardless of the shuffle method, both variants result in only minor fluctuations in node classification accuracy, indicating that prompt-based GNNs with the subsequent fine-tuning are robust to input permutation and maintain stable performance.
Table 13: Permutation sensitivity of prompt-based GNNs.
      Variant       Cora       Citeseer
      shuffle nodes       90.64 $\pm$1.12       83.79 $\pm$ 0.89
      shuffle tokens       90.55 $\pm$1.03       84.28 $\pm$1.06
Appendix H Case Study
H.1 Textual Representations in Prompt-based GNNs
In this part, we select two representative cases in citation networks and demonstrate their layer-by-layer refinement of our prompt-based GNN. Specifically, we provide the textual representations at each layer, including Round 0 as initial features. More importantly, we also collect its one-hop nodes and highlighted the source and relevant information below. From our empirical studies, we have the following observations. Overall, it is evident that the verbose textual representations are progressively refined to concise textual presentations. Meanwhile, the core semantics become increasingly clear throughout the process, until a short sequence composed of several natural language tokens at the last round. Furthermore, we notice that we effectively incorporate the key ideas of some neighboring nodes, as reflected in the refined textual representations after each aggregation-update operation. More specifically, the neighboring information is naturally fused to the central node within the text space, boosting their connections and improving the expressiveness of the central node. This study further illustrates the explicit message passing through our proposed prompt-based GNN in the text space, which captures the graph semantics and structures simultaneously.
H.1.1 Case A
Raw Text Attributes. Paper Title: Evaluating Neural Network Predictors by Bootstrapping. Abstract: We present a new method, inspired by the bootstrap, whose goal it is to determine the quality and reliability of a neural network predictor. Our method leads to more robust forecasting along with a large amount of statistical information on forecast performance that we exploit.We exhibit the method in the context of multi-variate time series prediction on financial data from the New York Stock Exchange. It turns out that the variation due to different resamplings (i.e., splits between training, cross-validation, and test sets) is significantly larger than the variation due to different network conditions (such as architecture and initial weights). Furthermore, this method allows us to forecast a probability distribution, as opposed to the traditional case of just a single value at each time step. We demonstrate this on a strictly held-out test set that includes the 1987 stock market crash. We also compare the performance of the class of neural networks to identically bootstrapped linear models.
Round 0: Textual Feature Initialization. This method evaluates neural network predictors using a bootstrap-inspired approach, improving robustness in NYSE financial data forecasting. It predicts probability distributions rather than single values, surpassing resampling variations. Compared to bootstrapped linear models, it enhances accuracy, as demonstrated during the 1987 stock market crash.
Round 1: 1st Layer Representations. This bootstrap method improves neural network evaluations in NYSE forecasting, comparable to traditional error estimation methods. It clears and simplifies data structures, enhancing accuracy and predicting probability distribution.
/***"comparable to traditional error estimation methods" comes from the paper: A comparison of some error estimates for neural network models. ***/
/***"It clears and simplifies data structures" comes from the paper: The Observer-Observation Dilemma in Neuro-Forecasting: Reliable Models From Unreliable Data Through CLEARNING. ***/
Round 2: 2nd Layer Representations. This bootstrap-inspired method refines neural network evaluations for NYSE forecasting, highlighting local error bars from nonlinear regression analysis, improving accuracy with data simplification techniques.
/***"local error bars from nonlinear regression analysis" comes from the paper: Learning Local Error Bars for Nonlinear Regression. ***/
Round 3: 3rd Layer Representations. Bootstrap method refines neural forecasts, integrates error bars, enhances accuracy.
H.1.2 Case B
Raw Text Attributes. Paper Title: Insulin dependent diabetes in Asians. Abstract: Type 1 diabetes is said to be extremely rare in children in India, where diabetes treated with insulin may be due to chronic pancreatic disease or malnutrition. To see whether typical type 1 diabetes occurred in Asian children in the United Kingdom, all known Asian children with diabetes in industrial West Yorkshire were ascertained. A total of 17 such children were studied; of these, seven were from three multiplex families and two fathers from these families had diabetes. All children were ketosis prone and developed diabetes while resident in the UK. There were significant increases in HLA-B8 and HLA-DR3 and increases in HLA-DR4 and HLA-DR3/DR4, while HLA-B15 was absent. Islet cell antibodies, either IgG or complement fixing, were present in four of 18 subjects tested, all of whom had disease of short duration. The prevalence of type 1 diabetes in Asian children aged 15 years or less in West Yorkshire was 36/100,000, assuming complete ascertainment. It is concluded that typical type 1 diabetes may occur in Asian children and this condition may be more common in families who have migrated to the UK.
Round 0: Textual Feature Initialization. The study examined type 1 diabetes prevalence in 17 Asian children in West Yorkshire, UK, focusing on family history. Seven were from multiplex families, and all were ketosis-prone. Significant increases in HLA-B8, HLA-DR3, and HLA-DR4 were found, with no HLA-B15. Islet cell antibodies were present in four of 18 subjects. The diabetes prevalence was 36/100,000, suggesting it may be more common in migrant families.
Round 1: 1st Layer Representations. The study examined type 1 diabetes in Asian children in West Yorkshire, UK. Environmental factors and migration may raise incidence, unlike the low incidence of insulin-dependent diabetes in Karachi. Increases in HLA-B8, HLA-DR3, and HLA-DR4 were found.
/***"Environmental factors and migration may raise incidence" comes from the paper: Evidence for an environmental effect in the aetiology of insulin dependent diabetes in a transmigratory population. ***/
/***"low incidence of insulin-dependent diabetes in Karachi" comes from the paper: Incidence of insulin dependent diabetes mellitus in Karachi, Pakistan. ***/
Round 2: 2nd Layer Representations. The study examined type 1 diabetes in Asian children in West Yorkshire, UK. Migration may raise incidence, unlike the low incidence in Karachi. Increases in HLA markers were found, similar to North and South Indian diabetics.
/***"Migration may raise incidence" comes from the paper: HLA-DR antigen frequencies in a North Indian type I diabetic population. ***/
/***"North and South Indian diabetics" comes from two papers: HLA-DR antigen frequencies in a North Indian type I diabetic population and HLA, complement C2, C4, properdin factor B and glyoxalase types in South Indian diabetics. ***/
Round 3: 3rd Layer Representations. Type 1, diabetes. Migration and HLA markers linked to increased diabetes incidence.
H.2 Language-based IDs vs. Key Words
To provide further insights, we leverage external information to validate the superiority of PromptGFM in capturing the core semantics of nodes. Specifically, we extract key words of papers from citation datasets (Cora, Citeseer, Obgn-arxiv, and PubMed) and compare them with their language-based IDs in our universal graph vocabulary.
Table 14 summarizes the key words and language-based IDs of selected papers, along with their titles and URLs for reference. Overall, it is evident that there are strong semantic relevance between the language-based IDs and keywords. For example, regarding the paper titled Distributed Protocols at the Rescue for Trustworthy Online Voting, the key words have appeared within its language-based ID, suggesting that PromptGFM has effectively captured its core semantics through our prompt-based GNN. In addition, in Committees providing EJR can be computed efficiently, where the title is less indicative of the content, the language-based ID still aligns perfectly with the corresponding key words, such as efficient computation and rules. This finding demonstrates that PromptGFM not only effectively captures the core idea without relying on the title, but also filters relevant semantics from neighboring nodes to enhance its own representations. Overall, our language-based IDs accurately capture and extend the semantics of the nodes, making them well-suited to form a universal graph vocabulary.
Table 14: The comparison between language-based IDs generated by prompt-based GNNs as part of the graph vocabulary and keywords extracted from the original papers in citation datasets. To prevent data leakage, we exclude these keywords from model training. The observed similarity confirms that our prompt-based GNN effectively captures the essential semantics of nodes. Furthermore, our language-based IDs encode richer information than keywords by preserving high-order signals.
Paper Language-based ID Key Words URL
Modular Verification of Interrupt-Driven Software Modular verification of interrupt-driven software using abstract interpretation Software, Abstract Interpretation, Feasibility Verification arXiv:1709.10078
Parsimonious Data: How a single Facebook like predicts voting behaviour in multiparty systems Predicting voting behavior using Facebook likes in multiparty systems Facebook Likes, Voter Intention, Machine Learning, Multiparty System arXiv:1704.01143
A Fast Noniterative Algorithm for Compressive Sensing Using Binary Measurement Matrices Fast noniterative algorithm for compressive sensing with binary matrices Compressive Sensing, Deterministic Methods arXiv:1708.03608
Optimization of Battery Energy Storage to Improve Power System Oscillation Damping Battery storage optimization improves power system oscillation damping Battery Energy Storage System, Oscillation Damping arXiv:1811.10213
Neural Variational Hybrid Collaborative Filtering Neural Variational Hybrid Collaborative Filtering improves recommendation performance Collaborative Filtering, VAE, Recommendation System arXiv:1810.05376v6
Interpretable Neural Networks for Predicting Mortality Risk using Multi-modal Electronic Health Records Predicting mortality risk using interpretable multi-modal neural network Mortality Risk Prediction, Clinical Data arXiv:1901.08125
A New Approach to Distributed Hypothesis Testing and Non-Bayesian Learning: Improved Learning Rate and Byzantine-Resilience Distributed hypothesis testing with Byzantine resilience using Bayesian update Bayesian Learning, Byzantine Resilience arXiv:1907.03588
Accurate and Efficient Hyperbolic Tangent Activation Function on FPGA using the DCT Interpolation Filter Efficient hyperbolic tangent activation function using DCTIF Hyperbolic Tangent, Activation Function arXiv:1609.07750
Distributed Protocols at the Rescue for Trustworthy Online Voting Trustworthy online voting with distributed blockchain protocols Distributed Voting, Distributed Protocols arXiv:1705.04480
Committees providing EJR can be computed efficiently Efficient computation of approval-based multi-winner voting rules Approval-Based Voting, Multi-Winner Elections arXiv:1704.00356
Creatism: A deep-learning photographer capable of creating professional work Creatism: deep learning system for artistic photography creation Creatism, Evaluation of Photographic Quality, Deep Learning arXiv:1707.03491
Relation of familial patterns of coronary heart disease, stroke, and diabetes to subclinical atherosclerosis: the multi-ethnic study of atherosclerosis Family history beyond early-onset heart disease impacts atherosclerosis Family History, Coronary Heart Disease, Stroke doi.org/10.1097/GIM.0b0
13e31818e639b
Glycemic index, glycemic load, and risk of type 2 diabetes Benefits of low-GI diet in type 2 diabetes Diabetes, Prevention doi.org/10.1093/ajcn/76/1.
274S
Decreased insulin responsiveness of glucose uptake in cultured human skeletal muscle cells from insulin-resistant nondiabetic relatives of type 2 diabetic families Inherited defects contribute to insulin resistance in diabetes Insulin Resistance, Inherited Factors doi.org/10.2337/diabetes.
49.7.1169
Quantitative histopathological studies of the extramural coronary arteries from Type 2 (non-insulin-dependent) diabetic patients Histopathological study of coronary arteries in diabetic patients Histopathology, Diabetes Mellitus doi.org/10.1007/BF0027
4798
Metabolic control and diet in Finnish diabetic adolescents Factors influencing metabolic control in diabetic adolescents Diabetes Mellitus, Adolescent doi.org/10.1111/j.1651-2227.1992.tb12212.x
H.3 Impact Statement
We highlight the following concrete cases to better demonstrate the broad applicability of our approach.
$\bullet$ In recommender systems, different scenarios involve distinct graph structures and scenario-specific semantics. Instead of training scenario-specific models, our foundation model could act as a general-purpose recommendation backbone that unifies multiple scenarios in text space via a co-training paradigm. This enables LLMs to capture rich and transferable semantic relationships across scenarios (e.g., user preferences, item affinities, contextual behaviors). Specifically, (a) our foundation model alleviates cross-dataset inconsistencies and facilitates rapid adaptation to new scenarios and tasks. (b) it is particularly effective in cold-start issues and long-tail recommendation challenges, where existing methods often struggle due to rigid architectures and heavy dependence on labeled data.
$\bullet$ In biomedical knowledge graphs, entities (e.g., genes, proteins, drugs, diseases) and their complex relations (e.g., inhibition, expression, association) form semantically rich graphs. Our approach achieves cross-resource knowledge integration (e.g., DrugBank, CTD, DisGeNET) and supports inductive reasoning across biomedical subdomains, enabling fast adaptation to new tasks and unseen relation types with minimal supervision. This is particularly valuable in biomedical settings, where knowledge resources are often fragmented, relation annotations are sparse, and entity semantics vary significantly across datasets. In this context, our approach is well suited for low-resource scenarios such as rare disease research, novel drug discovery, and cross-database knowledge transfer.