### v3

# LLM as GNN: Graph Vocabulary Learning for Text-Attributed Graph Foundation Models

Abstract.
Text-Attributed Graphs (TAGs), where each node is associated with text descriptions, are ubiquitous in real-world scenarios. They typically exhibit distinctive structure and domain-specific knowledge, motivating the development of Graph Foundation Models (GFM) that generalize across diverse graphs and tasks. Despite large efforts to integrate Large Language Models (LLMs) and Graph Neural Networks (GNNs) for TAGs, most methods suffer from decoupled architectures with two-stage alignment, limiting their synergistic potential. Even worse, existing methods assign out-of-vocabulary (OOV) tokens to graph nodes, leading to graph-specific semantics, token explosion, and incompatibility with task-oriented prompt templates, which hinders cross-graph and cross-task transferability. To address these challenges, we propose PromptGFM, a versatile GFM for TAGs grounded in graph vocabulary learning. PromptGFM comprises two key components: (a) graph understanding module, which explicitly prompts LLMs to replicate the finest GNN workflow within the text space, facilitating seamless GNN-LLM integration and elegant graph-text alignment; (b) graph inference module, which establishes a language-based graph vocabulary ensuring expressiveness, transferability, and scalability, enabling readable instructions for joint LLM fine-tuning across graphs and tasks. Extensive experiments demonstrate our superiority, especially effective transferability across diverse graphs and tasks. The code is available at https://github.com/agiresearch/PromptGFM.
Graph-Text Alignment, Foundation Models, Transfer Learning
copyright: nonecopyright: none
1. Introduction
Graphs, characterized by their non-Euclidean structures and rich domain-specific knowledge, serve as fundamental representations of complex relational data. A wide range of real-world graphs are further enriched with node-level textual information, forming Text-Attributed Graphs (TAGs), such as citation networks  (Eto, 2019; Hu et al., 2020; Buneman et al., 2021), social networks  (Kempe et al., 2003; Myers et al., 2014), and molecular graphs  (Wieder et al., 2020; Jin et al., 2024). However, existing approaches often rely on task- or dataset-specific training paradigms, hindering their transferability. To overcome this limitation, we seek to develop a Graph Foundation Model (GFM) capable of generalizing across different graphs and tasks for TAGs  (Mao et al., 2024; Xia et al., 2024).
Figure 1. Overview of three GNN-LLM integration paradigms for graph-text alignment: (a) GNN for LLM and (b) LLM for GNN use decoupled architectures in the embedding space, while (c) our work functions LLM as GNN in the text space.
In TAGs, Graph Neural Networks (GNNs) and Large Language Models (LLMs) are employed to process different modalities, as illustrated in Figure 1. (a) GNN for LLM. GNNs generate structure-aware node embeddings that complement original textual embeddings, improving LLM inference  (Tang et al., 2024; Chai et al., 2023; Liu et al., 2024b). (b) LLM for GNN. LLMs extract additional semantic features or labels from textual data, serving as supervision signals for GNN training  (Chen et al., 2024b; Liu et al., 2024a; Zhu et al., 2024). However, existing two-stage alignment pipeline, where GNNs and LLMs are only loosely coupled, struggles to fully exploit their complementary strengths, often leading to suboptimal graph-text alignment.
Recently, a noteworthy trend has emerged toward using LLM as GNN, where graph verbalizers convert graph data into heuristic or code-like prompts, enabling LLMs to understand graph semantics and structure simultaneously  (Ye et al., 2024; Wang et al., 2024a; Chen et al., 2024c). However, we argue that these approaches do not qualify as true GNNs, as they omit the core message-passing mechanism. As illustrated in Figure 2(a), a traditional GNN layer includes neighbor sampling, aggregation-update, and optimization  (Kipf and Welling, 2017; Velickovic et al., 2017). By stacking such layers, structure-less embeddings gradually evolve into structure-rich embeddings that encode higher-order dependencies. Without these essential components, a key challenge arises: Can we leverage LLMs to faithfully replicate GNNs to jointly capture graph semantics and structure?
Meanwhile, existing works intuitively treat each graph node as an out-of-vocabulary (OOV) token, resulting in graph-specific semantics and uncontrolled vocabulary explosion  (Tang et al., 2024; Ye et al., 2024). Worse still, due to vocabulary mismatches, ID-based node embeddings and language-based token embeddings reside in different feature spaces, causing semantic misalignment during LLM inference. This incompatibility undermines both transferability and scalability of graph-specific knowledge across diverse graphs and tasks. To enable knowledge transfer, an urgent question emerges: Can we replace OOV tokens with compatible and universal node representations to build a versatile GFM?
Figure 2. LLM-driven replication of the GNN workflow. We achieve fine-grained alignment between traditional embedding-based GNN and our prompt-based GNN.
A versatile GFM should be grounded in graph vocabulary learning  (Cai, 2024; Mao et al., 2024; Wang et al., 2024b) with the following desired properties: (1) Expressiveness. It encapsulates both semantic and structural information across diverse graphs. (2) Transferability. Every node in any graph should be representable by a composition of fundamental units within the vocabulary. (3) Scalability. The vocabulary should be inclusive to accommodate unseen nodes, even beyond existing graphs. Since natural language is inherently an expressive medium built upon meaningful and transferable tokens (Raffel et al., 2020; Radford et al., 2021; Palo et al., 2023; Wang et al., 2024b), we are inspired to establish a universal graph vocabulary in the text space for node representations. Therefore, we propose PromptGFM as a GFM for TAGs, which is consisted of two components.
Graph Understanding Module. To function LLMs as GNNs, we initialize node features with textual attributes and prompt LLMs to explicitly replicate the fine-grained GNN workflow in text space. As shown in Figure 2, we encode local structure via one-hop neighbor descriptions and design prompts that guide a flexible aggregation-update mechanism. To support optimization, we incorporate heuristic prompts that reflect contrastive learning at each layer. Finally, iterative LLM calls simulate the message passing over the whole graph, progressively refining verbose textual features into concise yet informative textual representations. Unlike traditional GNNs that output numerical embeddings, our prompt-based GNN produces structured text encodings that preserve both node semantics and high-order connections. Consequently, LLMs can operate as GNNs, and GNNs can be interpreted as LLMs, unlocking the full potential of GNN-LLM integration and empowering elegant graph-text alignment.
Graph Inference Module. Having captured semantics and structure through prompt-based GNNs, we decouple the resulting textual representations to establish a graph vocabulary, where each node is mapped to a finite sequence of language-based tokens, essentially as language-based IDs. Since these IDs are directly drawn from the LLM’s native vocabulary, they inherently preserve semantic fidelity and avoid OOV issues. As a result, this vocabulary is universally interpretable, transferable across graphs, and scalable to unseen nodes, thereby resolving semantic irrelevance and graph-text incompatibility. Afterward, language-based IDs can be seamlessly integrated to generate readable and coherent natural-language prompts. To support broad generalization, we adopt a multi-instruction fine-tuning framework that gathers diverse instructions from multiple graphs and tasks, facilitating cross-graph and cross-task knowledge transfer. In conclusion, this graph vocabulary not only eliminates graph-text incompatibility but also paves the way for general GFMs.
The main contributions of this work are summarized as follows:
$\bullet$ We present PromptGFM, a graph foundation model for TAGs built on graph vocabulary learning.
$\bullet$ We introduce a novel LLM-as-GNN paradigm with a text-space prompt-based GNN, enabling seamless GNN–LLM integration and effective graph–text alignment.
$\bullet$ We construct a language-based graph vocabulary as the cornerstone of the GFM, ensuring graph-text semantic fidelity and eliminating the OOV issue.
$\bullet$ We perform extensive experiments to validate the superiority of PromptGFM and systematically analyze its cross-graph and cross-task transferability, together with its adaptability to diverse pre-training tasks, establishing it as a general and versatile GFM.
Figure 3. The pipeline of PromptGFM. (a) Graph Understanding Module: For arbitrary TAGs from different domains, prompt-based GNN replicates traditional embedding-based GNN workflow in the text space, generating compact node representations. (b) Graph Inference Module: We establish a unified graph vocabulary and extract language-based IDs to generate massive pure-language prompts, enabling LLM fine-tuning across graphs and tasks.
2. Related Works
GNN-LLM Integration. LLMs have unlocked unprecedented potential for graph machine learning, driving efforts to integrate GNNs and LLMs for modeling TAGs. (a) GNN for LLM. As structure tokenizers, GNNs and graph transformers embed graph topology into textual representations, enhancing semantical understanding in LLMs (Tang et al., 2024; Chai et al., 2023; Liu et al., 2024b). However, effectively coordinating GNN-LLM architectures and co-training remains a challenge. (b) LLM for GNN. LLMs assist GNNs by generating node-level and edge-level labels to address data sparsity issues  (Chen et al., 2024b; Guo et al., 2024; Xia et al., 2024). Meanwhile, the other line directs LLMs to produce additional features or explanations to overcome semantic deficiencies  (Liu et al., 2024a; Zhu et al., 2024; He et al., 2024). However, their reliance on LLM-generated content inevitably introduces noise, impacting performance. (c) LLM as GNN. This paradigm directly operates LLMs as GNNs by designing structure verbalizers to convert graph data into heuristic or code-like prompts for LLM inference  (Chen et al., 2024c; Ye et al., 2024; Wang et al., 2024a). Nevertheless, they fail to capture high-order connections due to the lack of an intrinsic GNN mechanism. Overall, the current decoupled approaches rely on two-stage alignment, failing to fully exploit the strengths of both models. This limitation motivates us to propose a new paradigm where LLMs inherently function as GNNs, maximizing their synergistic potential.
Graph-Text Alignment in Embedding Space. Current approaches to graph-text alignment primarily operate in the embedding space. One approach uses graph encoders as prefixes, mapping graph-aware embeddings to language-based embeddings for LLM fine-tuning in a shared space  (Huang et al., 2023, 2024). Other works adopt a two-tower architecture, leveraging contrastive learning  (Li et al., 2023a; Brannon et al., 2023; Tang et al., 2024), iterative training  (Zhao et al., 2023a; Zhu et al., 2024), or knowledge distillation  (Mavromatis et al., 2023) to align distinct representations. However, these methods encounter a persistent modality gap, limiting transferability and scalability across graphs. Since TAGs contain textual information, we advocate for shifting graph-text alignment to the text space.
Graphs Foundation Models. A GFM aims to achieve transferability across different datasets and tasks, where the key challenge lies in finding a graph vocabulary that identifies transferable units to encode invariance on graphs  (Liu et al., 2023; Mao et al., 2024). Previous works rely on domain-specific vocabularies: GraphGPT  (Tang et al., 2024) assumes unique IDs to nodes and creates a dataset-specific vocabulary, while MoleBERT  (Xia et al., 2023) defines a molecular graph vocabulary by converting atomic properties into chemically meaningful codes. Despite their success, they lack in-context learning and cross-domain transferability. Recently, while some studies have investigated understanding and inferring graphs in natural language  (Fatemi et al., 2024; Wang et al., 2023; Zhao et al., 2023b; Liu et al., 2024a), none of them have attempted to establish a language-based graph vocabulary. To address this, we introduce an expressive graph vocabulary to exploit the inherent transferability of native natural language, advancing the development of a more versatile GFM.
3. Preliminaries
Text-Attributed Graphs (TAGs). A TAG is defined as ${G}=\left(V,E,X\right)$, where $V$ is the set of nodes, $E$ is the set of edges, and $X$ is the set of textual node attributes. Each node $v_{i}\in{V}$ is associated with a textual description ${X}_{i}=\left(x_{i}^{1},x_{i}^{2},\ldots,x_{i}^{{n_{i}}}\right)$, where each $x_{i}^{k}\in\mathcal{X},k=1,\dots,{n_{i}}$. Here, $\mathcal{X}$ denotes the natural language dictionary.
Graph Neural Networks. GNNs have emerged as state-of-the-art models in graph machine learning, fundamentally built upon the message-passing mechanism. In practice, a GNN selects neighboring nodes to the target node, aggregates their representations to capture local structure, and updates the target node’s representation. Mathematically, for a node $v_{i}$, the $l$-th layer of a general GNN is defined as:
(1) $\mathcal{N}_{i}^{(l)}$ $=f_{\text{sample}}\left(\mathcal{N}_{i}\right),$
$\mathbf{m}_{i}^{(l)}$ $=f_{\text{agg}}\left({\left\{\mathbf{h}_{j}^{(l-1)},v_{j}\in\mathcal{N}_{i}^{(l)}\right\}}\right),$
$\mathbf{h}_{i}^{(l)}$ $=f_{\text{update}}\left(\mathbf{h}_{i}^{(l-1)},\mathbf{m}_{i}^{(l)}\right)$
where $\mathbf{h}_{i}^{(l)}$ is the node embedding of $v_{i}$ at the $l$-th layer. $\mathcal{N}_{i}$ denotes full neighbors and $\mathcal{N}_{i}^{(l)}$ is sampled neighbors at the $l$-th layer. To capture high-order connections, we stack $L$ layers and derive final embeddings as $\mathbf{h}_{i}=f_{\text{pooling}}\left(\mathbf{h}_{i}^{(1)},...,\mathbf{h}_{i}^{(L)}\right)$  (Grattarola et al., 2024). For optimization, a contrastive loss with negative sampling is commonly used in unsupervised graph learning  (Hamilton et al., 2017; Velickovic et al., 2019), expressed as:
(2) $\begin{split}\ell=f_{\text{loss}}\left(f_{\text{sim}}\left(\mathbf{h}_{i},\mathbf{h}_{j}\right),f_{\text{sim}}\left(\mathbf{h}_{i},\mathbf{h}_{k}\right)\right),\end{split}$
where $v_{j}\in\mathcal{N}_{i}^{(l)}$ is a positive sample and $v_{k}\notin\mathcal{N}_{i}$ is a negative one. $f_{\text{sim}}\left(\cdot\right)$ measures the similarity between two nodes. $f_{\text{loss}}\left(\cdot\right)$ enforces contrastive learning by increasing similarity for connected nodes and reducing it for unconnected ones.
Figure 4. An instance of graph inference module in link prediction, where language-based IDs are indexed from the graph vocabulary to generate readable instructions with task-oriented templates. A multi-instruction fine-tuning framework is employed to unify graphs and tasks and learn transferable knowledge for GFMs.
4. Methodology
This section presents the pipeline of our proposed PromptGFM, which consists of a graph understanding module and a graph inference module, as illustrated in Figure 3.
4.1. Graph Understanding Module
The graph understanding module aims to generate expressive representations for each node within the graph to support downstream inference. The main challenge lies in effectively aligning semantic and structural information, where LLMs excel in textual understanding and GNNs in structural modeling  (Li et al., 2023b; Ren et al., 2024). To bridge the gap in GNN-LLM integration, we propose prompt-based GNNs that operate LLMs as GNNs by prompting LLMs to replicate general GNN workflow within the text space.
Our priority is to design appropriate prompts that enable LLMs to replicate GNNs. This requires considering three essential factors: (a) Graph Descriptions. How can we effectively describe the node semantics and graph structure to LLMs? (b) Graph Structure. How can we simulate message passing to encode structural dependencies? (c) Graph Semantics. How can we distill essential semantics into concise but informative textual representations?
As illustrated in Figure 2, we implement a fine-grained replication of a general GNN workflow, i.e. prompt-based GNN. We begin by initializing node representations by summarizing their textual attributes, analogous to the lookup layer in conventional GNNs. To reduce computational overhead, we then sample one-hop neighbor information as follows:
(3) $\left\{X_{j}^{(l-1)},\left\{v_{j}\right\}\subset\mathcal{N}_{i}\right\}$ $\leftarrow\text{Prompt}_{\text{sample}}\left(X^{(l-1)},\mathcal{N}_{i}\right),$
where $X^{(l-1)}=\left\{X_{0}^{(l-1)},X_{1}^{(l-1)},\dots,X_{|V|-1}^{(l-1)}\right\}$ denotes textual representations of all nodes from the previous layer, and $\left\{X_{j}^{(l-1)}\right\}$ corresponds to the selected neighbors $\mathcal{N}_{i}$. Then, we directly prompt LLMs to perform the core message aggregation-update process for each node, formulated as:
(4) $X_{i}^{(l)}$ $\leftarrow\text{Prompt}_{\text{agg-update}}\left(\left\{X_{j}^{(l-1)},\left\{v_{j}\right\}\in\mathcal{N}_{i}^{(l)}\right\},X_{i}^{(l-1)}\right),$
where we seek flexible message passing mechanism without relying on predefined operators such as mean and weighted aggregators. A discussion on handling prompt length constraints is provided in Appendix C.1. In terms of optimization, we refer to unsupervised graph learning with a prompt-driven contrastive loss, which encourages similarity among neighboring nodes while pushing apart distant ones. As a task-agnostic approach to graph representation learning, contrastive optimization naturally aligns with the generality requirements of GFMs. Instead of traditional negative sampling, which becomes redundant in our setting, we prompt LLMs to intuitively and implicitly steer the contrastive objective. Details of prompt construction are provided in Appendix B.
After repeating $L$ rounds, we obtain the final textual representation as $T_{i}=X_{i}^{(L)}$. These representations encapsulate both semantic and structural information, effectively solving the aforementioned issues: (a) Graph Description. One-hop neighbor descriptions are equivalent to an adjacency matrix, representing the entire graph structure. (b) Graph Structure. To capture high-order relationships, we iteratively invoke LLMs with the same prompts, feeding each round’s output into the next. (c) Graph Semantics. Simultaneously, we explicitly instruct LLMs to produce concise yet expressive textual representations for each node, gradually refining them for denser and richer semantics. In summary, our fine-grained replication exemplifies the potential of using LLMs as GNNs, fostering seamless GNN-LLM integration and elegant graph-text alignment.
4.2. Graph Inference Module
The graph inference module seeks to unify diverse graphs and tasks and acquire transferable knowledge through multi-instruction fine-tuning with LLMs. Existing methods, which represent nodes as OOV tokens, creates semantic misalignment between ID-based and language-based embeddings in task-oriented prompts due to distinct vocabularies, and substantially limits transferability across graphs and tasks. To overcome this challenge, we introduce a novel language-based graph vocabulary that bridges this incompatibility, enabling the generation of large-scale, coherent, and readable instructions to support effective LLM inference.
4.2.1. Graph Vocabulary.
Learning a transferable graph vocabulary, whose fundamental units can represent each node, is critical for building effective GFMs. Its utility hinges on three key criteria: expressiveness, transferability, and scalability. Since each node has been equipped with a textual representation in the graph understanding module that captures its core semantics and local structure, we intuitively construct a universal language-based graph vocabulary that matches LLM’s native vocabulary. In our graph vocabulary, each node is represented by a finite sequence of language tokens, known as a language-based ID, defined as:
(5) $\mathcal{F}:V\rightarrow T,$
where each graph node $v_{i}$ maps a sequence ${T}_{i}=\left(t_{i}^{1},t_{i}^{2},\ldots,t_{i}^{m_{i}}\right)$, and $t_{i}^{k}\in\mathcal{X}$ and $\mathcal{X}$ is a dictionary of general natural language tokens. Collectively, the language-based IDs for all nodes in $V$ are denoted as $T=\left\{T_{0},T_{1},\dots,T_{|V|-1}\right\}$. Along this line, the structured token sequence of a graph node is analogous to a word in standard lexicons, ensuring semantic fidelity and eliminating OOV issue.
Our vocabulary satisfies all expected criteria: (a) Expressiveness. Language-based IDs preserve rich semantic and structural information from open-world settings. (b) Transferability. Like human lexicons, the vocabulary is grounded in natural language, providing a shared foundation that ensures cross-graph interpretability and transferability. (c) Scalability. Any node, whether previously seen or not, can dynamically generate its language-based ID, keeping compatibility with existing nodes while preventing token explosion. We analyze its coverage and uniqueness in Section 5.6.
4.2.2. Multi-Instruction Fine-Tuning.
As shown in Figure 4, we employ a multi-instruction fine-tuning framework to incorporate diverse graphs and tasks (Chung et al., 2024; Wei et al., 2022). We index nodes from the vocabulary and embed their language-based IDs into task-oriented templates to construct completed instructions $\mathcal{T}$:
(6) $\mathcal{T}\leftarrow\text{Prompt}_{\text{template}}\left(T,G\mid\mathcal{F}\right),$
which include language-based IDs of the central nodes, along with its corresponding local structure from $G$ depending on specific graph-centric tasks (e.g., node classification or link prediction). As these instructions are composed entirely of natural language tokens, they are fully readable and coherent. All graph-centric tasks are then reformulated in a text-to-text format, allowing the LLM to be fine-tuned within a unified instruction learning paradigm (Mishra et al., 2022). Given a target sequence $Y$, the loss function can be computed as:
(7) $\mathcal{L}$ $=-\sum_{j=1}^{|Y|}\log\text{Pr}\left(y_{j}\mid\mathcal{T},Y_{<j}\right),$
where $\text{Pr}\left(y_{j}\mid\mathcal{T},Y_{<j}\right)$ denotes the probability of the $j$-th token $Y_{j}$ in the output sequence $Y$, conditioned on the instruction $\mathcal{T}$ and all previous tokens $Y_{<j}=(y_{1},y_{2},\dots,y_{j-1})$. To support broad generalization, we fine-tune LLM backbones such as T5 (Raffel et al., 2020), FLAN (Wei et al., 2022), and Llama (Touvron et al., 2023) to co-train a unified model across diverse graphs and tasks. As a result, the model acquires open-world knowledge and generalizes effectively to previously unseen graphs and tasks.
4.2.3. Constrained Decoding via Prefix Tree Search.
To mitigate LLM hallucination, we introduce a constrained decoding strategy based on a prefix tree search to regulate LLM outputs  (Cao et al., 2021; Tan et al., 2024). Taking generative link prediction as an example, we craft a prefix tree from language-based IDs of all candidate nodes, where each tree node corresponds to a natural language token and each root-to-leaf path uniquely identifies a valid node ID. During autoregressive generation, token selection is constrained to a valid path within the prefix tree, ensuring that predictions align with actual graph nodes and eliminating hallucinations. This effectiveness can be attributed to the discrete nature of language-based IDs, further highlighting the flexibility of our proposed graph vocabulary.
4.3. Efficiency Analysis.
4.3.1. Graph Understanding Module.
Although PromptGFM incurs an initial API expense during the graph understanding phase, this cost is a one-time overhead and remains manageable with modern infrastructure. Each graph is processed only once to generate language-based IDs, which are cached and reused across the inference pipeline. Consequently, when new graphs are introduced, existing IDs do not need to be regenerated via repeated API calls; instead, additional LLM fine-tuning over these natural language tokens is sufficient to capture updated features.
4.3.2. Graph Inference Module.
With IDs cached, both fine-tuning and decoding operate entirely on tokens of these IDs. At inference time, the complexity per decoding step is $O(L)$, where $L$ is the average ID length. Prefix tree search constrains the candidate space to graph-specific vocabulary size ($N$), the overall decoding complexity becomes $O(N\times L)$. These properties make the computational overhead manageable, even in the large-scale graph setting.
Table 1. Evaluation results ($\%$) on node classification accuracy ($\uparrow$). We highlight the methods with the first and second best performances, and report the statistical significance over the strongest baseline using a two-tailed $t$-test with $p<0.05$.
Method Cora Citeseer PubMed Ogbn-arxiv History Photo WikiCS
MLP 62.29 64.42 62.88 62.07 64.62 61.21 68.41
GCN 82.47 76.11 77.36 66.15 81.93 78.58 76.33
GAT 82.92 77.30 74.36 65.29 82.85 82.38 78.21
SAGE 83.69 73.17 83.22 68.78 82.12 80.06 79.56
RevGNN 86.90 77.34 82.16 70.43 83.04 83.24 81.22
SGFormer 82.36 73.76 78.92 63.44 78.98 80.12 76.56
NodeFormer 81.55 72.98 76.49 73.21 79.60 78.51 75.47
BERT 79.02 72.83 76.74 71.90 72.97 68.82 77.98
Sentence-BERT 78.09 72.12 75.48 77.24 74.10 69.02 77.72
Llama3 78.17 74.68 77.46 77.65 76.49 74.42 79.94
OFA 75.72 71.58 75.26 74.68 81.43 84.46 78.02
LLaGA 81.25 68.80 86.54 76.05 82.55 85.34 80.74
ENGINE 91.48 78.46 90.24 76.02 82.46 83.75 81.56
GraphPrompter 80.26 73.61 94.80 75.61 79.42 80.04 80.98
PromptGFM (Flan-T5) 91.72 84.49 92.83 80.58 82.33 85.41 81.49
PromptGFM (Llama3-8B) 92.42 85.32 94.65 83.78 86.72 86.61 84.66
pp-value 4e-3 2.73e-8 3.21e-7 4.48e-10 4.85e-7 3.62e-2 9.93e-4
5. Experiments
In this section, we conduct extensive experiments to address the following research questions (RQs): RQ1: How does PromptGFM perform on node classification and link prediction tasks? RQ2: How to transfer across diverse graphs and tasks as a general-purpose GFM? RQ3: How does PromptGFM accommodate more pretraining tasks? RQ4: What factors influence the replication of prompt-based GNN? RQ5: What is the computational efficiency in graph understanding and inference modules?
5.1. Experimental Setup
5.1.1. Datasets.
We present following datasets from three domains: (1) Computer Science: Cora (McCallum et al., 2000), Citeseer (Giles et al., 1998), Ogbn-arxiv (Hu et al., 2020), and WikiCS  (Mernyei and Cangea, 2020); (2) E-commerce: Photo and History (Ni et al., 2019; Yan et al., 2023); (3) Biomedical: PubMed (Sen et al., 2008). Details can be found in Appendix A.1.
5.1.2. Baselines.
We make comprehensive comparisons with existing methods in four categories: (1) Graph-agnostic methods. We consider the MLP model without graph structure. (2) GNN-based methods. We employ four fundamental GNN models: GCN  (Kipf and Welling, 2017), GAT (Velickovic et al., 2017), GraphSAGE (Hamilton et al., 2017), and ReVGNN (Li et al., 2021). We also explore SGFormer (Wu et al., 2023) and NodeFormer (Wu et al., 2022), which leverage Transformer architectures to model graph data. (3) LM-only methods. We utilize BERT (Devlin et al., 2019), Sentence-BERT  (Reimers and Gurevych, 2019), and Llama3-8B (Touvron et al., 2023) to address downstream tasks. (4) GNN-LLM Integration methods. Following the proposed taxonomy, GraphPrompter  (Liu et al., 2024b) is an instance of using GNNs to enhance LLMs. OFA  (Liu et al., 2024a) and ENGINE  (Zhu et al., 2024) are examples of leveraging LLMs for GNNs. Besides, we incorporate LLaGA  (Chen et al., 2024c) as an attempt of implementing LLM as GNN. Details can be found in Appendix A.2.
5.1.3. Reproduction Settings.
We conduct all experiments in PyTorch on four NVIDIA RTX A6000 GPUs. The graph understanding module uses OpenAI’s GPT-4o mini while we fine-tune a Flan-T5 or Llama3-8B in the graph inference module. We employ 10-fold cross-validation and compute average results with standard deviations. For evaluation, we use accuracy and Macro-F1 (M-F1) for node classification, and accuracy and HR@1 for link prediction. To align with the TAG setting, BERT is used for node initialization across all applicable baselines. More details are provided in Appendix A.3.
Table 2. Accuracy ($\%$) on discriminative link prediction.
Method Discriminative: Accuracy ($\uparrow$)
Cora Citeseer Obgn-arxiv PubMed
GCN 77.15 78.72 80.89 77.36
GAT 70.44 77.17 76.25 74.36
SAGE 85.31 87.15 80.76 83.22
GraphPrompter 90.10 91.67 73.21 80.49
Ours (Flan-T5) 90.57 92.03 81.12 87.64
Ours (Llama3) 91.68 93.46 84.27 88.12
Table 3. HR@1 ($\%$) on generative link prediction.
Method Generative: HR@1 ($\uparrow$)
Cora Citeseer PubMed
GCN 5.95 6.82 0.51
GAT 2.22 3.59 0.28
SAGE 6.59 8.73 0.45
PromptGFM 8.21 8.90 1.21
5.2. Performance Comparison (RQ1)
5.2.1. Main Results.
We train each model independently from scratch on a single graph and compare their performance. Tables 1 and 2 summarize the results for node classification and link prediction accuracy, respectively. PromptGFM achieves substantial improvements over state-of-the-art models, leading to the following observations. (a) Graph-based models consistently outperform graph-agnostic ones, highlighting the importance of structural information. (b) Our method surpasses OFA, ENGINE, and GraphPrompter on node classification, suggesting the limitations of their decoupled architectures in both GNN-for-LLM and LLM-for-GNN paradigms. (c) PromptGFM also surpasses LLaGA, which relies on templates to implicitly understand graph semantics and structures. These results indicate that heuristic prompting alone is insufficient to capture high-order dependencies without a true GNN mechanism. In contrast, PromptGFM effectively replicates message passing through a prompt-based GNN, demonstrating the feasibility of using LLMs as GNNs and establishing a new paradigm for graph-text alignment. We provide further analysis in Section 5.5.
5.2.2. Generative Link Prediction.
Following a transductive setting, we partition the graph by links, construct an input graph from the training edges, and predict missing connections for test nodes, treating existing nodes as candidate targets. From Table 3, PromptGFM consistently outperforms traditional GNN models on generative link prediction. Existing GNN-LLM approaches, however, fail in this task: their reliance on OOV token embeddings prevents direct mapping from LLM outputs to specific nodes, often leading to hallucinated predictions. In contrast, our method represents nodes as finite token sequences and employs constrained decoding through prefix tree search, which regulates LLM outputs. This further underscores the significance of our graph vocabulary and demonstrates the flexibility of PromptGFM in generative scenarios.
Table 4. Intra-domain cross-graph transferability to Citeseer within the computer science domain.
Source Pre-training Co-training
Cora Arxiv Acc($\uparrow$) M-F1($\uparrow$) Acc($\uparrow$) M-F1($\uparrow$)
27.64 17.10 84.49 82.31
51.63 45.10 84.96 83.22
60.34 54.81 85.45 83.91
61.25 55.66 86.77 84.24
Table 5. Inter-domain cross-graph transferability from e-commerce to biomedical domain.
Source Pre-training Co-training
History Photo Acc($\uparrow$) M-F1($\uparrow$) Acc($\uparrow$) M-F1($\uparrow$)
39.12 39.84 90.67 91.82
46.24 47.41 88.52 89.39
44.21 44.93 85.34 85.57
45.18 45.65 86.22 87.03
Table 6. Inter-domain cross-graph transferability from computer science to biomedical domain.
Source Pre-training Co-training
Cora Citeseer Arxiv Acc($\uparrow$) M-F1($\uparrow$) Acc($\uparrow$) M-F1($\uparrow$)
39.12 39.84 90.67 91.82
51.76 52.84 86.34 87.29
40.12 42.38 85.21 86.12
60.21 62.02 82.17 86.01
50.17 51.74 84.39 81.94
57.28 59.71 81.32 83.14
55.34 57.11 82.13 83.10
53.07 54.90 80.43 80.79
5.3. Cross-graph and Cross-task Transfer (RQ2)
We investigate cross-graph transferability in both intra-domain and inter-domain scenarios. Following  (Chen et al., 2024a), we explore two settings: (a) Pre-training, where models are evaluated on entirely unseen datasets without prior exposure during training, and (b) Co-training, where a portion of the target dataset is available and jointly used with auxiliary datasets for training. These two dimensions yield four distinct transfer paradigms, enabling a comprehensive assessment of cross-graph transferability as follows.
5.3.1. Intra-domain Cross-graph Transferability.
As illustrated in Table 4, all settings surpass direct inference using an off-the-shelf LLM. Notably, incorporating Arxiv yields greater gains than Cora, suggesting that larger and richer datasets enhance transferability more effectively. In the pre-training setting, substantial improvements over direct inference demonstrate strong zero-shot transferability within the domain. A similar, albeit smaller, trend is observed in the co-training setting, further validating the effectiveness of intra-domain knowledge transfer. These findings underscore the potential to collect more graph data and build more knowledgeable and transferable GFMs.
5.3.2. Inter-domain Cross-graph Transferability.
We report cross-domain transfer results from different domains to the biomedical domain in Tables 5 and 6, respectively. In the pre-training setting, while incorporating a single external graph always enhances performance, adding multiple source graphs may lead to negative transfer. More critically, in the co-training setting under supervised learning, performance consistently deteriorates as more cross-domain data is introduced. This decline mainly stems from domain shifts, including structural differences, semantic drift in text, and mismatched label spaces  (Chen et al., 2024a; Zhao et al., 2024). In contrast, PromptGFM adopts a unified language-based vocabulary and pure language interface, which enables natural cross-domain transferability and avoids token collision in principle. While this remains a broader challenge for the community, we provide additional discussion and supporting experiments in Appendix C.2.
Table 7. Cross-task generalization between node classification (NC) and link prediction LP. ST-X refers to supervised training on the task X from scratch.
Task Setting Cora Citeseer PubMed
NC zero-shot 18.54 27.64 39.12
[…]
ST-NC 91.72 84.49 92.83
LP zero-shot 67.28 69.22 73.59
[…]
ST-LP 91.68 93.46 88.12
5.3.3. Cross-task Transferability.
[…]
5.4. Diverse Pretraining Tasks (RQ3)
A robust foundation model should support a rich set of pre-training tasks, as task diversity is expected to enhance generalization on downstream applications. To this end, we extend beyond the NC and LP tasks by introducing two additional pre-training objectives: (a) Node Discrimination (ND): Given three nodes [a], [b], [c], exactly one belongs to a different category. Which one is it?. and (b) Link Discrimination (LD): Given a central node [X], which candidate node [a], [b], [c] is not connected to [X]?, where [x] denotes the language-based ID of a node. Besides NC and LP, we incorporate ND and LD individually or jointly into multi-task fine-tuning, and evaluate their effects on NC performance under otherwise identical settings. As shown in Table 8, we observe that adding either ND or LD improves both accuracy and M-F1, while combining the two yields the best performance. These results validate that expanding the pre-training task set leads to richer representations. More importantly, they demonstrate that PromptGFM does not merely memorize task patterns, but leverages its unified language-based interface to generalize across tasks and exhibit emergent capabilities, a defining property of foundation models.
Table 8. The impact of incorporating Node Discrimination (+ND) and Link Discrimination (+LD) as pre-training tasks on NC performance.
Dataset +ND +LD Accuracy (%) Macro-F1 (%)
Citeseer 84.49 82.31
85.67 ($\uparrow$1.18) 83.22 ($\uparrow$0.91)
84.73 ($\uparrow$0.24) 82.65 ($\uparrow$0.34)
86.02 ($\uparrow$1.53) 83.77 ($\uparrow$1.46)
PubMed 90.67 91.82
92.53 ($\uparrow$1.86) 92.84 ($\uparrow$1.02)
91.12 ($\uparrow$0.45) 92.03 ($\uparrow$0.21)
93.06 ($\uparrow$2.39) 93.27 ($\uparrow$1.45)
Table 9. Permutation sensitivity of prompt-based GNNs.
Variant Cora Citeseer
shuffle nodes 90.64 $\pm$1.12 83.79 $\pm$ 0.89
shuffle tokens 90.55 $\pm$1.03 84.28 $\pm$1.06
Table 10. Layer-wise mean cosine similarity of node embeddings on Cora, with margin reflecting the gap between connected and disconnected pairs.
Connected Pairs Disconnected Pairs Margin
Layer 1 0.171 0.036 0.135
Layer 2 0.223 0.032 0.191
Layer 3 0.308 0.034 0.274
5.5. Prompt-based GNN Analysis (RQ4)
5.5.1. Permutation Sensitivity
We evaluate the sensitivity of our prompt-based GNNs (using Flan-T5) to input order by evaluating two variants: (a) shuffle nodes, which randomizes the order of neighbors, and (b) shuffle tokens, which injects position-specific tokens to nodes and shuffle them instead. As illustrated in Table 9, regardless of the shuffle method, both variants result in only minor fluctuations in node classification accuracy, indicating that prompt-based GNNs with the subsequent fine-tuning are robust to input permutation and maintain manageable and stable performance.
Table 11. Quantitative graph vocabulary analysis.
Metric Cora Citeseer PubMed Ogbn-arxiv
Coverage (%) 100.00 100.00 100.00 100.00
Entropy 9.1116 9.4561 8.4643 10.7464
Metric History Photo WikiCS ALL
Coverage (%) 100.00 100.00 100.00 100.00
Entropy 10.3752 11.3272 9.3825 12.2761
Figure 5. Ablation studies on node classification.
Figure 6. Impact of layer numbers.
5.5.2. Representation Discrimination over Layers
To validate that our language-based graph vocabulary effectively preserves structural proximity, we conduct an analysis by encoding the final language-based IDs with a pretrained Sentence-BERT encoder and computing the cosine similarity between node pairs. The results on the Cora dataset are presented in Table 10. We observe that connected nodes exhibit a significantly higher mean similarity than disconnected nodes ($p<10^{-5}$). Furthermore, the margin between the two groups grows steadily as the number of GNN layers increases, indicating stronger structural discrimination with deeper propagation. This trend highlights that our prompt-based GNN captures meaningful semantics with stacking aggregation-update layers, thereby demonstrating its theoretical equivalence in representational power to traditional embedding-based GNNs.
5.5.3. Performance over GNN Layers.
We investigate the impact of layer depth in our prompt-based GNN. As shown in Figure 6, increasing the number of layers initially improves results by capturing broader context and higher-order dependencies. However, beyond a certain depth, performance plateaus or declines due to over-smoothing, a phenomenon also observed in traditional GNNs. The optimal depth is achieved with three layers on Cora and two layers on Citeseer. These results suggest that textual representations can be propagated across the graph in much the same way as numerical embeddings, effectively encoding both semantic and structural information. We further provide a case study on layer-wise representations in Appendix C.3.
5.6. Exploration Studies
5.6.1. Ablation Studies.
We construct three ablated variants to assess the contribution of each component: (a) w/o understanding, which removes the prompt-based GNN; (b) w/o inference, which eliminates multi-instruction fine-tuning; and (c) w/o both, which directly prompts LLMs using raw textual attributes. Figure 6 reports the node classification accuracy, where our full model consistently outperforms all variants. Specifically, the performance drop in w/o understanding highlights the importance of replicating GNN workflow to preserve both semantic and structural signals. The sharp decline in w/o inference further emphasizes the necessity of LLM fine-tuning for incorporating domain-specific knowledge. Finally, w/o both yields the worst results, underscoring the complementarity between the graph understanding and inference modules.
5.6.2. Vocabulary Analysis.
We evaluate the quality of our graph vocabulary using two metrics in Table 11. (a) Coverage. We compute the proportion of nodes successfully matched to IDs. Since the graph vocabulary is grounded in LLM’s native language, any node’s semantics can, in principle, be expressed through a combination of natural language tokens. Thus, every node is assigned a valid ID, confirming full coverage. (b) Uniqueness. We measure the entropy of token distributions across nodes within both individual and merged graphs. Higher entropy indicates more diverse and fine-grained token usage, indicating stronger uniqueness and representational capacity. Notably, the merged vocabulary (“ALL”) achieves the highest entropy (12.2761), showing that integrating multiple graphs improves vocabulary expressiveness and discriminability, which reinforces the necessity of building GFMs.
5.6.3. Efficiency Analysis.
We evaluate the computational and monetary cost of PromptGFM, as summarized in Table 12.In the graph understanding module, the API or token cost grows with the graph size, but this constitutes a one-time, graph-specific overhead that can be cached and reused throughout the inference pipeline (denoted as ”n/a”). During inference, as more graphs are introduced and the overall vocabulary expands, we observe that time and energy consumption increase in an approximately linear manner. These results align with our discussion in Section 4.3 and and demonstrate PromptGFM’s efficiency as a scalable GFM.
Table 12. Efficiency analysis in terms of API cost (USD), number of tokens (M), inference time (min), and power consumption (kWh).
Source Understanding Inference
Cora Citeseer Photo API Cost # Tokens Time Power
0.188 1.25 16.47 0.07
0.189 1.29 21.41 0.08
1.213 8.08 83.67 0.34
n/a n/a 35.31 0.16
n/a n/a 117.56 1.04
6. Conclusion
In this work, we present PromptGFM, a GFM for TAGs built on graph vocabulary learning. With GNN replication within the text space, we decouple refined textual node representations and establish a unified graph vocabulary. This vocabulary endows compatibility and scalability, facilitating effective LLM fine-tuning with readable instructions for enhanced transferability. Experiments validate its superior performance and effective cross-graph and cross-task generalization. Our research reveal the potential of using LLM as GNN, opening new avenues for developing GFMs in TAGs.
Appendix A Experimental Setup
A.1. Data Descriptions
We use seven public benchmarking datasets from three domains to evaluate PromptGFM, and the statistics are illustrated in Table 13.
A.2. Baselines
We provide detailed information on the baseline models, categorized into: (1) Graph-agnostic methods, (2) GNN-based methods, (3) LM-only methods, and (4) GNN-LLM integration methods.
A.2.1. Graph-agnostic methods.
$\bullet$ MLP. This method adopts a multi-layer perceptron to learn low-dimensional embeddings for each node.
A.2.2. GNN-based methods.
$\bullet$ GCN  (Kipf and Welling, 2017). This model generalizes convolutional operations to graphs for semi-supervised learning.
$\bullet$ GAT  (Velickovic et al., 2017). This method integrates attention mechanisms to weigh neighbor features during aggregation.
$\bullet$ SAGE  (Hamilton et al., 2017). GraphSAGE is an inductive framework which enables generalization to unseen data.
$\bullet$ ReVGNN  (Li et al., 2021). This method includes a recurrent GNN that updates node representations as the graph evolves over time.
$\bullet$ SGFormer  (Wu et al., 2023). This work introduces a Transformer-based architecture that integrates spectral filters to capture both local and global graph structures.
$\bullet$ NodeFormer  (Wu et al., 2022). This framework presents a scalable graph Transformer that utilizes a randomized attention mechanism, enabling efficient training on large-scale graphs.
A.2.3. LM-only methods.
$\bullet$ BERT  (Devlin et al., 2019). This language model learns bidirectional context through pre-training with masked language modeling (MLM) and next sentence prediction (NSP).
$\bullet$ Sentence-BERT  (Reimers and Gurevych, 2019). This model is a modification of BERT that uses a siamese network to generate fixed-size sentence embeddings.
$\bullet$ LLaMA3  (Liu et al., 2024a). This is an open-weight, decoder-only Transformer with 8 billion parameters, which achieves strong performance on a wide range of benchmarks, particularly in reasoning, multilingual tasks, and instruction following.
A.2.4. GNN-LLM integration methods.
$\bullet$ LLaGA  (Chen et al., 2024c). This model integrates LLMs for graph data by transforming nodes into structure-aware sequences and mapping them into the token embedding space with a projector.
$\bullet$ OFA  (Liu et al., 2024a). This framework handles various graph classification tasks across different domains using a single model.
$\bullet$ GraphPrompter  (Liu et al., 2024b). This framework aligns graphs with LLMs via soft prompts, including a GNN to capture structure and an LLM to interpret node text, combining semantic and structural insights through prompt tuning.
$\bullet$ ENGINE  (Zhu et al., 2024). This paper presents a parameter- and memory-efficient fine-tuning method, which combines a tunable GNN-based side structure alongside each LLM layer.
A.3. Implementation Details
We provide additional details to facilitate reproducibility. In the graph understanding module, the number of layers in the prompt-based GNN is selected from $\{1,2,3,4\}$, and the first-order neighbor sampling ratio from $\{0.3,0.6,1.0\}$. In the graph inference module, we fine-tune the LLM with a learning rate of 3e-4 and a batch size of 4. To mitigate potential biases introduced by task-specific prompts, we designed a prompt pool for each task requirement and randomly selected prompts during instruction construction to enhance robustness. We employed a standard early-stopping strategy during training: if the performance metric on the validation set did not improve over a fixed number of consecutive epochs (determined based on the dataset), we halted training to prevent overfitting. For other hyperparameters of the compared methods, we referred to the original papers and carefully tuned them to suit each dataset.
Figure 7. Graph understanding module via prompt-based GNNs. We prompt LLMs to achieve fine-grained reproduction of traditional GNN workflow, refining verbose textual representations into concise yet meaningful ones. In the prompt, neighbor sampling (see Equation 3) is in blue, the aggregation-update mechanism (see Equation 4) in blue, and the optimization in red.
Appendix B Prompt Design
This section provides templates of prompts, including prompt-based GNNs and task-oriented prompts.
Prompt for node initialization. The title of the paper is ¡the title of the paper¿, the abstract of the paper is ¡the abstract of the paper¿. Please summarize the paper. /*** This prompt varies depending on the dataset. The present instance is for citation datasets only. ***/
Prompt for each layer of GNN replication. Given the central node ¡l-th round textual representation of the central node¿. The selected one-hop neighbors are [¡ l-th round of node #1¿, ¡l-th round node #2¿, … ,¡l-th round node #N¿]. Please aggregate neighbor nodes and update a concise yet meaningful representation for the central node. Note connected nodes should share similar semantics and vice versa.
Prompt for node classification. ¡central node language-based ID¿ has 1-hop connections with […, ¡1-hop neighbor language-based IDs¿, …], and it also has 2-hop connections with […, ¡2-hop neighbor language-based IDs¿, …]. Which category should ¡central node language-based ID¿ be classified as?
Prompt for discriminative link prediction. ¡central node language-based ID ¿ has 1-hop connections with […, ¡1-hop neighbor language-based IDs¿, …], and it also has 2-hop connections with […, ¡2-hop neighbor language-based IDs¿, …]. Among ¡central node language-based ID¿ and ¡negative sampling node language-based ID¿, which node will be connected to ¡central node language-based ID¿ ?
Prompt for generative link prediction. ¡central node language-based ID¿ has 1-hop connections with […, ¡1-hop neighbor language-based IDs¿, …], and it also has 2-hop connections with […, ¡2-hop neighbor language-based IDs¿, …]. Which node should be connected to ¡central node language-based ID¿?
Table 14. Analysis of High-Degree Nodes and Average Connectivity
Metric Cora Citeseer PubMed Ogbn-arxiv History Photo WikiCS
Avg. Degree 4.01 2.84 13.77 36.90 18.07 12.11 4.50
High-degree (¿300) Nodes 0 0 0 0.20% 0.25% 0.19% 0
Appendix C Further Exploration
C.1. Prompt Length Limitation Discussion
In dense TAGs, high-degree nodes present a challenge for prompt-based GNNs, as their long prompts may exceed LLM input limits and hinder efficiency. To address this issue, we examine strategies that balance efficiency, scalability, and representational fidelity.
C.1.1. Neighbor Sampling in Scalable GNNs.
Following standard practices, we adopt a neighbor sampling strategy in each GNN layer to reduce prompt length and enhance robustness. As shown in Table 15, sampling 60% of neighbors yields the best overall performance across datasets. Meanwhile, limiting the maximum sampled neighbors is a common practice in scalable GNNs. For example, GraphSAGE samples 25 and 10 neighbors at the first and second hops, while PinSAGE selects from $\{10,20,50\}$. Following this rationale, we explore $\{10,20,30,40\}$ and empirically set 20 as the cap, balancing LLM input constraints with sufficient structural coverage.
C.1.2. Token Capacity of LLMs.
Our implementation leverages GPT-4o mini, which supports up to 128k input tokens per call. This capacity allows a single prompt to encode information from over 300 neighbors, given that each neighbor contributes about 400 tokens in the first propagation round. As shown in Table 14, only a small fraction of nodes (¡0.25%) in a few datasets surpass this capacity. For these rare high-degree nodes, our neighbor sampling and truncation mechanism can bound prompt length, ensuring broad scalability to dense graphs without information loss.
C.1.3. Prompt Design.
Prompt-based GNNs guide the LLM to aggregate and update neighbor information through structured prompts, producing increasingly compact yet expressive representations. Across layers, the average length shrinks from  400 to  10 tokens as semantic and structural information are progressively distilled. Since each node is processed independently and depends only on its local degree, prompt size remains bounded regardless of graph scale, ensuring both efficiency and scalability to large graphs.
Table 15. The performance with varying neighbor sampling ratios acorss three datasets.
Sampling Ratio 100% 60% 30%
Cora 81.36 89.35 88.36
Citeseer 79.34 82.23 81.13
PubMed 90.23 91.12 89.34
Table 16. The effectiveness of adaptive prompting that alleviates inter-domain negative transfer.
Source Pre-training Co-training
Acc ($\uparrow$) M-F1 ($\uparrow$) Acc ($\uparrow$) M-F1 ($\uparrow$)
History Photo w/o w/ w/o w/ w/o w/ w/o w/
39.12 40.23 39.84 40.13 90.67 90.82 91.82 91.93
46.24 47.44 47.41 47.98 88.52 90.27 89.39 90.22
44.21 45.11 44.93 46.02 85.34 88.13 85.57 88.47
45.18 47.90 45.65 47.26 86.22 89.02 87.03 89.16
C.2. Negative Transfer Discussion
C.2.1. Why does performance drop?
Cross-domain negative transfer mainly stems from domain shifts, including structure heterogeneity, semantic shift in text, and label space mismatch (especially for node classification), rather than hard vocabulary conflicts. In contrast, PromptGFM adopts a unified language-based vocabulary and pure language interface, which enables natural cross-domain transferability and avoids token collision in principle.
C.2.2. Adaptive Prompting Exploration.
[…]
C.3. Case Study
We visualize the layer-by-layer refinement of node representations and identify which neighbor nodes contribute to each stage. Textual representations evolve from verbose descriptions into compact, high-density token sequences (8-10 tokens). After the first layer, nodes absorb neighbor semantics (e.g., ”error estimation”); after the second round, they integrate broader two-hop context (e.g., ”nonlinear regression analysis”). These observations demonstrate the explicit message passing behavior of the prompt-based GNN in text space, effectively capturing both graph semantics and structural dependencies.
Raw Text Attributes. Paper Title: Evaluating Neural Network Predictors by Bootstrapping. Abstract: We present a new method, inspired by the bootstrap, whose goal it is to determine the quality and reliability of a neural network predictor. Our method leads to more robust forecasting along with a large amount of statistical information on forecast performance that we exploit.We exhibit the method in the context of multi-variate time series prediction on financial data from the New York Stock Exchange. It turns out that the variation due to different resamplings (i.e., splits between training, cross-validation, and test sets) is significantly larger than the variation due to different network conditions (such as architecture and initial weights). Furthermore, this method allows us to forecast a probability distribution, as opposed to the traditional case of just a single value at each time step. We demonstrate this on a strictly held-out test set that includes the 1987 stock market crash. We also compare the performance of the class of neural networks to identically bootstrapped linear models.
Round 0: Textual Feature Initialization. This method evaluates neural network predictors using a bootstrap-inspired approach, improving robustness in NYSE financial data forecasting. It predicts probability distributions rather than single values, surpassing resampling variations. Compared to bootstrapped linear models, it enhances accuracy, as demonstrated during the 1987 stock market crash.
Round 1: 1st Layer Representations. This bootstrap method improves neural network evaluations in NYSE forecasting, comparable to traditional error estimation methods [A]. It clears and simplifies data structures [B], enhancing accuracy and predicting probability distribution.
/***[A] comes from the paper: A comparison of some error estimates for neural network models. ***/
/***[B] comes from the paper: The Observer-Observation Dilemma in Neuro-Forecasting: Reliable Models From Unreliable Data Through Clearning. ***/
Round 2: 2nd Layer Representations. This bootstrap-inspired method refines neural network evaluations for NYSE forecasting, highlighting local error bars from nonlinear regression analysis [C], improving accuracy with data simplification techniques.
/***[C] comes from the paper: Learning Local Error Bars for Nonlinear Regression. ***/
Round 3: 3rd Layer Representations. Bootstrap method refines neural forecasts, integrates error bars, enhances accuracy.